### v1

# Video Soundtrack Generation by Aligning Emotions and Temporal Boundaries

Abstract
We introduce EMSYNC, a video-based symbolic music generation model that aligns music with a video’s emotional content and temporal boundaries. It follows a two-stage framework, where a pretrained video emotion classifier extracts emotional features, and a conditional music generator produces MIDI sequences guided by both emotional and temporal cues. We introduce boundary offsets, a novel temporal conditioning mechanism that enables the model to anticipate and align musical chords with scene cuts. Unlike existing models, our approach retains event-based encoding, ensuring fine-grained timing control and expressive musical nuances. We also propose a mapping scheme to bridge the video emotion classifier, which produces discrete emotion categories, with the emotion-conditioned MIDI generator, which operates on continuous-valued valence-arousal inputs. In subjective listening tests, EMSYNC outperforms state-of-the-art models across all subjective metrics, for music theory-aware participants as well as the general listeners.
1 Introduction
The online distribution of user-generated multimedia content is expanding at an exponential rate thanks to affordable, high-quality recording equipment and video editing software Falkowski-Gilski and Uhl (2020). A key challenge in content creation is providing suitable soundtracks to enhance viewer engagement Buhler (2018). However, unauthorized use of commercially published music infringes copyright, preventing monetization for creators on platforms like YouTube. Alternatives such as purchasing music, hiring composers, or searching for royalty-free tracks are often costly, time-consuming, or fail to ensure proper synchronization with video content. Automatic video-based music generation offers a promising solution to this problem Di et al. (2021); Kang et al. (2024); Zhuo et al. (2023).
Existing approaches that generate music as audio waveforms lack editability Lam et al. (2024); Copet et al. (2024). This method compresses all stages of music production—composition, performance, editing, mixing, and mastering—into a single process, limiting creative control for professionals. By contrast, generating music in a symbolic format like MIDI (Musical Instrument Digital Interface) offers greater flexibility. MIDI functions as a digital music score, encoding instrument names, note pitches, durations, and velocities. Professionals can edit compositions, synthesize audio using digital audio workstations (DAWs), or record and refine performances.
This work focuses on generating music in MIDI format from arbitrary videos using deep neural networks (DNNs). We use “MIDI” and “music” interchangeably, as MIDI is our exclusive format. A major challenge in training video-to-MIDI DNNs is the absence of large-scale paired video-MIDI datasets. Existing datasets are either domain-specific, such as pianist hand videos Koepke et al. (2020), or contain limited samples, around 1k Di et al. (2021); Zhuo et al. (2023). To address this, we develop a model that generates music for any video type, leveraging the Lakh MIDI Dataset—the largest available MIDI dataset with 176,581 samples—ensuring diverse and high-quality outputs Raffel (2016).
Since the Lakh MIDI dataset lacks corresponding videos, we adopt a two-stage approach: extracting video features relevant to music generation and using them as conditioning inputs. We name our model EMSYNC, as it aligns music with video by matching their emotions, and synchronizing their temporal boundaries. While our method is applicable to any selection of temporal boundary and emotion representation, we define musical boundaries using long-duration chords, video boundaries using scene cuts, and represent emotions with the valence-arousal model Russell (1980). Specifically, we extract scene cut locations from the input video and guide the music generator to produce long-duration chords near those locations that are rhythmically and harmonically compatible with the rest of the generated music. We also use a pretrained video emotion classifier to estimate discrete emotion probabilities Sulun et al. (2024a), map them to valence-arousal values, and condition our music generator accordingly.
Temporal conditioning in MIDI-generating models presents unique challenges. While deep transformer models can handle time-based data, such as videos, their sequence dimension correlates linearly with time due to a fixed frame rate Liu et al. (2022); Arnab et al. (2021). In contrast, MIDI is typically processed using an event-based representation, where sequence and time dimensions are not linearly correlated Oore et al. (2020). In event-based MIDI encoding, two primary token types exist: “note” and “time shift”. Note tokens represent pitch, while time shift tokens advance the time axis, capturing both note durations and silences. Each time shift token specifies a time increment, forming a one-dimensional sequence where the position in the sequence does not directly correspond to the position in time. The key advantage of event-based encoding is the absence of a fixed time grid, allowing for subtle, expressive timing variations that reflect human musicianship. Unlike state-of-the-art video-based music generators that rely on fixed, coarse time grids Di et al. (2021); Kang et al. (2024), we introduce a novel temporal conditioning method that preserves event-based encoding, enabling fine-grained temporal control.
We also tackle the challenge of using emotions as an intermediary to link music generation to video. Sulun et. al Sulun et al. (2022) previously assigned valence-arousal emotion labels to samples in the Lakh MIDI dataset Raffel (2016), creating a training dataset two orders of magnitude larger than existing emotion-labeled MIDI datasets Panda et al. (2013); Ferreira and Whitehead (2019); Hung et al. (2021). However, the largest emotion-labeled video datasets primarily provide categorical labels Xu et al. (2016); Sun et al. (2020). To bridge this gap, we employ the VEMOCLAP model (Video EMOtion Classifier using Pretrained features) Sulun et al. (2024a), trained on the Ekman-$6$ dataset Xu et al. (2016). Since VEMOCLAP outputs probabilities of categorical emotions, we map these to valence-arousal values using prior user study results on emotion Russell and Mehrabian (1977).
We compare our method to state-of-the-art video-based MIDI generators through subjective listening tests, where it outperforms all competing methods across all metrics for both music theory-aware participants and general listeners while also being the fastest. Our key contributions include the following:
  • We develop a video-based MIDI generator that surpasses state-of-the-art models across all subjective metrics and participant demographics.
  • We introduce boundary offsets for temporal conditioning in transformers, ensuring precise alignment between the input video and the output MIDI.
  • We leverage insights from psychological studies to map discrete emotion categories to continuous valence-arousal values, enabling an emotional connection between the input video and the generated music.
  • We design our model architecture to modify only the inputs to the transformer body, allowing future users to fine-tune it for other tasks and data structures seamlessly.
2 Related work
In this section, we introduce the video emotion classifier and the emotion-based MIDI generator used in our work. We then review existing video-based MIDI generation methods and discuss their evaluation approaches.
2.1 Video emotion classification
The VEMOCLAP model exploits publicly available pretrained models for a multimodal video analysis and uses the resulting pretrained features to classify the emotions of arbitrary videos Sulun et al. (2024a, b). It employs pretrained models for automatic speech recognition (ASR), optical character recognition (OCR), facial expression classification, audio classification, and image understanding. Textual features obtained from ASR and OCR are further processed using a text sentiment classifier. Cross-attention layers Vaswani et al. (2017) integrate and process these multimodal features, and a linear layer produces the final emotion probabilities. The model is trained on the Ekman-6 dataset Xu et al. (2016), consisting of 1637 videos labeled with the six basic emotions derived from the original work of Ekman: anger, disgust, fear, joy, sadness, and surprise Ekman (1971).
2.2 Emotion-based MIDI generation
Sulun et al. Sulun et al. (2022) collected valence and arousal values for songs in the Lakh Pianoroll Dataset Dong et al. (2018) and trained a conditional transformer using the resulting labeled dataset. To integrate continuous-valued valence and arousal features with discrete musical note tokens, they project valence and arousal into a vector space using separate linear layers. Musical input tokens are projected into vectors of the same dimensionality using an embedding layer. These vectors are then concatenated with the projected valence and arousal vectors along the sequence dimension and fed into the transformer body with relative global attention Huang et al. (2018).
2.3 Video-based MIDI generation
While our method applies to arbitrary videos, several works focus on generating symbolic music for specific types of videos, such as those featuring human movements like dancing or instrumental performances. The Foley Music model Gan et al. (2020) generates MIDI from videos of musicians by processing body keypoint movements using a Graph Convolutional Network Kipf and Welling (2017) and a Transformer Vaswani et al. (2017). Similarly, Koepke et al. Koepke et al. (2020) and Su et al. Su et al. (2020) utilize use deep neural networks with residual connections to generate symbolic music from videos of finger movements on piano keyboards. Due to their specialized nature, these approaches rely on datasets containing video-MIDI pairs, though these datasets typically contain fewer than 1k samples Zhao et al. (2018); Li et al. (2018); Koepke et al. (2020); Su et al. (2020). The RhythmicNet model employs a multi-stage process to generate music from dance videos by predicting beats and style, generating a drum track, and subsequently creating multitrack music Su et al. (2021).
Some studies explore the more general task of generating symbolic music for arbitrary videos. The most similar to our approach, the Controllable Music Transformer (CMT), generates music based on video features such as motion speed, motion saliency, and timing Di et al. (2021). CMT employs the Lakh Pianoroll Dataset Dong et al. (2018), the same dataset we use, and processes music using an extended Compound Word representation, where each token encodes type, beat/bar marking, note strength, note density, instrument, pitch, and duration Hsiao et al. (2021). While this representation reduces sequence length, it significantly increases input dimensionality. Consequently, CMT uses a low temporal resolution of four ticks per beat, translating to a time resolution of 125 milliseconds for a song at 120 beats per minute. Furthermore, note events align precisely with beat subdivisions, whereas human musicians introduce expressive timing deviations. In contrast, our model utilizes an event-based representation with independent time shift tokens at an 8-millisecond resolution. This resolution is sufficiently fine to capture musical nuance while maintaining absolute time rather than rigid beat subdivisions Oore et al. (2020). Additionally, unlike CMT, our model integrates high-level emotional conditioning alongside low-level temporal features.
In a follow-up to CMT, Zhuo et al. introduced the Symbolic Music Videos dataset, containing video-MIDI pairs Zhuo et al. (2023). They sourced MIDI data from piano tutorial videos, automatically transcribing audio using the Onsets and Frames model, resulting in 1140 samples Hawthorne et al. (2018). Video features such as color histograms, RGB frame differences for motion, and high-level CLIP features Radford et al. (2021) were extracted to train three models sequentially: one generating chord sequences, another generating the melody, and a third generating the accompaniment.
The Video2Music model shares similarities with our approach by utilizing both low-level video features and high-level emotional conditioning but differs in application Kang et al. (2024). The authors compiled the MuVi-Sync dataset, consisting of 748 music videos labeled with musical attributes like note density, loudness, chords, and key. Their encoder-decoder transformer takes low- and high-level video features, along with user-provided primer chord and key, to generate chord sequences. These sequences are then arpeggiated using fixed patterns to create the final MIDI output. However, this method’s reliance on a fixed time grid omits the subtle expressive timing found in human performances. Additionally, its use of fixed arpeggiation patterns limits musical diversity, and requiring user-defined chord and key inputs restricts accessibility for non-musicians.
2.3.1 Evaluation methods
Evaluation of video-based music generators is non-trivial. For objective evaluation, most works measure the difference between generated and ground-truth samples using metrics such as Pitch Class Histogram Entropy, Grooving Pattern Similarity, and Structureness Indicator Wu and Yang (2020). When the dataset contains paired video and MIDI samples, this method is straightforward Gan et al. (2020); Zhuo et al. (2023). For unpaired video and MIDI samples, objectively evaluating a model that generates MIDI from arbitrary videos is not feasible.
The authors of the Controllable Music Transformer, using unpaired datasets, evaluate output MIDIs generated unconditionally, without input video Di et al. (2021). They empirically show that unconditionally generated MIDI samples exhibit metrics closer to the unpaired MIDI dataset compared to video-conditioned output samples. This occurs because constraining the model with video structure causes deviations from intrinsic MIDI structures, and the best way to match a specific MIDI dataset’s metrics is to train exclusively on that dataset. Subjective evaluation through user studies remains the most common, and often the only, viable method for assessing video-based MIDI generators Di et al. (2021); Kang et al. (2024); Zhuo et al. (2023); Gan et al. (2020); Su et al. (2021).
3 Methodology
In this section, we present our video-based symbolic music generator. We match the output music to the input video based on emotions and temporal boundaries. Since there are no datasets where videos are paired with symbolic music, we employ a two-stage approach, i.e., use independent modules for video analysis and conditional music generation, and then combine them. Our pipeline is shown in Figure 1. We first highlight our music generator and how it is conditioned on emotions and temporal boundaries. We then describe how we incorporate temporal and emotional video features into our conditional music generator.
3.1 Conditional music generator
Our music generator is conditioned on emotions as valence-arousal values and temporal boundaries. Figure 2 illustrates the model, with the emotion-conditioning mechanism in the upper part and the boundary-conditioning mechanism in the lower part.
Figure 1: Video-based music generation pipeline. The text and the image next to the arrows demonstrate sample values.
3.1.1 Training dataset and preprocessing
We train our music generator on the Lakh Pianoroll Dataset (LPD) Dong et al. (2018), which contains 174,154 pianorolls derived from the Lakh MIDI Dataset Raffel (2016). We tokenize the pianorolls using an event-based symbolic music representation Oore et al. (2020). Specifically, an ON (note on) token marks the start of a note, and an OFF (note off) token marks its end. These tokens also encode pitch and instrument information. For example, a piano note with a MIDI pitch of 60 (C4) is denoted as PIANO_ON_60. Since a larger number of instruments increases vocabulary size, we use the Lakh Pianoroll Dataset-5 variant, where all instrument tracks are merged into five predefined categories: bass, drums, guitar, piano, and strings Dong et al. (2018). However, our method is adaptable to datasets with different instrument groupings.
TIMESHIFT tokens are used to move along the time axis, representing both note durations and the silences between them. Each token specifies a time increment in milliseconds. For example, an 800-millisecond shift is encoded as TIMESHIFT_800. We use a temporal resolution of 8 milliseconds with a maximum shift of 1000 milliseconds. Longer durations are represented using multiple consecutive TIMESHIFT tokens. We also use the START tokens to mark the beginning of the songs, the BAR tokens to indicate the musical bars, and the PAD tokens to standardize input sequence lengths in minibatches. We use the CHORD token to mark long-duration chords as temporal boundaries, as they often serve as anchor points in musical compositions Bharucha and Krumhansl (1983). However, this is a design choice, and our method can accommodate any selection of temporal boundaries.
For emotion-conditioned generation, we use the model architecture and valence-arousal labeled dataset of Sulun et al., detailed in Section 2.2, which includes 23,661 samples Sulun et al. (2022). To utilize the remaining unlabeled samples, we further modify the model to accept NaN (Not A Number) values for valence or arousal, enabling both conditional and unconditional generation. If valence or arousal is unspecified, we assign it a NaN value and use a learned vector in place of its projection vector.
Figure 2: Our music generator. Numbers underneath valence and arousal are sample values.
3.1.2 Temporal boundary matching
In addition to matching emotions, we also temporally synchronize the input video and output music by aligning their temporal boundaries. We define temporal boundaries as scene cut locations in video and chord locations in music. Since we employ a hybrid approach, we first train a music generator capable of incorporating boundary locations as input and producing music with chords near these boundaries. To achieve this, we label the chords in the Lakh Pianoroll Dataset. During training, we consider only guitar and piano chords with at least three simultaneous notes that last for a minimum of two beats. These chords are labeled by inserting a CHORD token before the first ON token of each chord. To allow the model to generate chords not only at video scene cut locations but also independently where musically appropriate, we randomly remove 20% of CHORD tokens during training. An example of a labeled chord is shown in the top row of Figure 3. In training, the temporal locations of chords from the ground-truth MIDI serve as input boundaries, whereas during video-based inference, we replace these boundaries with video scene cut locations.
Figure 3: Graphical illustration of boundaries. Top: In symbolic music, a chord with three or more simultaneous notes and a duration exceeding a set threshold defines a musical boundary and is used during training. Middle: In video, scene cuts serve as video boundaries and are used during inference. Bottom: Boundary offset represents the temporal distance to the next boundary. These figures are illustrative; offsets do not perfectly align with the music except at the boundary.
Chords are integral to a melody, providing harmonic and rhythmic support to surrounding notes, both preceding and following Bharucha and Krumhansl (1983). During inference, forcing a CHORD token into the sequence at a specific location may cause the chord to sound off-beat or overly abrupt. This occurs because the model, having no prior knowledge of the upcoming chord, may generate preceding notes that do not align naturally with it. To address this, we develop a method that enables the model to “anticipate” upcoming chords and generate preceding notes and time shifts accordingly. Additionally, we train the model to generate the CHORD token itself, ensuring rhythmic consistency in the generated music. To achieve this, we define boundary offsets for each input token, representing the remaining time until the next boundary. These offsets are capped at a maximum value and, since our music generator is autoregressive, they are computed based on future chords rather than past ones. Furthermore, we amplify the loss of the CHORD token by a factor of 10, as it influences multiple preceding chord notes and plays a critical role in structuring the generated music. Sample offsets are illustrated in the bottom row of Figure 3.
We process boundary offsets through a feed-forward network, generating boundary offset encodings, which are then concatenated with learned positional encodings Wang et al. (2020) along the feature dimension, as shown in the lower part of Figure 2. This architectural choice is motivated by several key factors. First, similar to the emotion-conditioning mechanism, we inject boundary offset encodings at the input level rather than within the transformer body, ensuring that the core model remains unchanged. This allows the model to process inputs with or without boundary offsets, enabling seamless fine-tuning by repeating the learned positional encodings along the feature dimension when boundary offsets are absent. Second, we avoid adding boundary offset encodings directly to learned positional encodings to maintain a distinction between the two. Finally, recent studies suggest that decoder-only transformers can implicitly learn positional encodings through their internal weights, even without explicitly adding them Haviv et al. (2022); Kazemnejad et al. (2024). Based on this insight, we halve the feature length of learned positional encodings and allocate the remaining feature space to boundary offset encodings. The resulting vector sequence consists of positional encodings augmented with boundary offset encodings. Following the standard transformer model, we add this sequence to the token embeddings Vaswani et al. (2017) before passing it to the transformer body with relative global attention Huang et al. (2018). The transformer’s input is represented as:
$\begin{split}\mathbf{X}_{\text{trans}}&=\text{Concat}_{s}\left(\text{FFN}_{v}(x_{v}),\text{FFN}_{a}(x_{a}),\text{Embedding}(\mathbf{x}_{t})\right)\\ &+\text{Concat}_{f}(\text{FFN}_{b}(\mathbf{b})+\mathbf{W_{pe}})\end{split}$
where $x_{v}$ and $x_{a}$ are the valence and arousal inputs, $\mathbf{x}_{t}$ is the input token sequence, $\mathbf{b}$ represents the input boundary offsets, and $\mathbf{W_{pe}}$ is the learned positional encoding. The feed-forward networks $\text{FFN}_{v}$, $\text{FFN}_{a}$, and $\text{FFN}_{b}$ process the valence, arousal, and boundary offset inputs, respectively. The operations $\text{Concat}_{s}$ and $\text{Concat}_{f}$ denote concatenation along the sequence and feature dimensions, respectively. Our music generator architecture is illustrated in Figure 2.
Algorithm 1 outlines the process of computing boundary offsets during inference, i.e., music generation. The model generates music autoregressively, producing one token per forward pass. By tracking the generated TIMESHIFT tokens, we maintain a time cursor and compute the boundary offset, i.e., the time remaining until the next boundary, for each generated token. If the model generates a CHORD token, we calculate the absolute difference between the time cursor and each input boundary. The boundary with a difference smaller than the predefined sensitivity threshold of 1 second is considered successfully generated. We then remove these boundaries from future offset calculations by replacing them with infinity. Next, we compute the boundary offset for the generated token, regardless of its type. This offset is determined as the distance to the next closest boundary, capped by a maximum offset value. For simplicity, Algorithm 1 is presented for a single sample, but in practice, this operation is performed in minibatches. During training, we preprocess the entire input sequence at once by constructing a time grid instead of a time cursor, allowing us to calculate boundary offsets for all tokens simultaneously. For clarity, the initial list of generated tokens is shown as empty; however, in practice, we begin with a START token.
The ON tokens that appear after a CHORD token and before the next TIMESHIFT token are considered notes of the generated chord. To make the generated chord more distinctive, we increase the velocity of these notes. Music generation continues until the time cursor reaches the duration of the video. After synthesizing the generated MIDI into an audio waveform, we apply a 3-second fade-out effect to prevent an abrupt ending and ensure the music matches the duration of the input video.
Algorithm 1 Creating boundary offsets during music generation in inference.
Input: List of input boundaries (in seconds), $\mathbf{b}$; video duration, $d$; valence, $x_{v}$; arousal $x_{a}$
Parameter: Sensitivity, $\xi$; maximum offset, $\delta_{max}$; generated boundary mask, $\mathbf{m_{b}}$; token type $t_{t}$; token value, $t_{v}$; time cursor, $c$; generation function including forward-pass and sampling, g
Output: List of generated tokens $\mathbf{t}$; list of boundary offsets, $\boldsymbol{\delta}$
  Let $c=0$; $\boldsymbol{\delta}=\left[\hskip 2.0pt\right]$; $\mathbf{t}=\left[\hskip 2.0pt\right]$.
  while $c<d$ do
     # generate token as (type, value):
     $(t_{t},t_{v})=g(\mathbf{t},\boldsymbol{\delta},x_{v},x_{a})$
     if $t_{t}$ == TIMESHIFT then
        $c=c+t_{v}$
     else if $t_{t}$ == CHORD then
        $\mathbf{m_{b}}=\left|t-\mathbf{b}\right|<\xi$.
        $\mathbf{b}\left[\mathbf{m_{b}}\right]=+\infty$.
     end if
     $\boldsymbol{\delta}$.append$(\max(\min(\mathbf{b}-c),\delta_{max}))$
     $\mathbf{t}$.append$((t_{t},t_{v}))$
  end while
3.2 Scene detection
During video-based inference, we first extract the video’s temporal boundaries, specifically its scene cut locations, using the FFmpeg software [^1] https://www.ffmpeg.org/. We then apply a difference filter to the extracted scene cuts and remove those occurring less than 4 seconds apart. The remaining timestamps serve as the input boundaries for the music generator.
3.3 Emotion matching
We use the VEMOCLAP emotion classifier to extract emotions from input videos as probability distributions over discrete emotion categories Sulun et al. (2024a). The music generator is conditioned on emotions represented as valence and arousal values ranging from -1 to 1, rather than discrete categories Sulun et al. (2022). Since the VEMOCLAP emotion classifier outputs discrete emotion categories, we first map them to the valence-arousal plane Russell (1980) before feeding them into our music generator.
Russell and Mehrabian conducted user studies to find the corresponding valence and arousal values of discrete emotion categories Russell and Mehrabian (1977). They presented their findings in a table that contains the mean and standard deviation values of valence and arousal for each categorical emotion. We extracted and used the values corresponding to Ekman’s six basic emotions, which are presented in Table 1 Ekman (1971).
Valence Arousal
Emotion Mean SD Mean SD
anger -0.51 0.20 -0.59 0.29
disgust -0.60 0.20 -0.35 0.41
fear -0.64 0.20 -0.60 0.32
joy -0.76 0.22 -0.48 0.26
sadness -0.63 0.23 -0.27 0.34
surprise -0.40 0.30 -0.67 0.27
Table 1: Means and standard deviations of valence and arousal values corresponding to Ekman’s categorical emotions Ekman (1971), obtained from the user studies of Russell and Mehrabian Russell and Mehrabian (1977).
Using the output probabilities of the video classifier along with the means and standard deviations of each emotion, we construct a mixture of Gaussian distributions. Users can either sample from the mixture—first selecting an emotion category based on the output probabilities and then sampling from the corresponding Gaussian distribution—or use the mean of the mixture, which is computed as the weighted average of the emotion category means, with weights determined by the output probabilities. Additionally, we introduce a parameter that represents the maximum absolute value of means across all emotion categories. Using this parameter, we compute a single scaling coefficient applied to all valence-arousal distributions. This approach allows users to adjust the conditioning of the music generator, choosing between a wider or narrower range of emotions.
Group 1 Group 2
MR MQ EM TM OM MR MQ EM TM OM RT
Video2Music 2.521 1.854 2.125 1.958 2.167 2.708 1.771 1.854 1.979 2.083 1.607
CMT 1.854 2.438 2.000 2.125 2.083 1.792 2.562 2.458 2.312 2.354 3.554
EMSYNC (Ours) 1.625 1.708 1.875 1.917 1.750 1.500 1.667 1.688 1.708 1.562 1.417
Table 2: Group 1 consists of participants with a self-reported understanding of music theory, while Group 2 includes those without it. We use the following abbreviations: MR (music richness), MQ (music quality), EM (emotion match), TM (timing match), OM (overall match), and RT (runtime). Runtime measures the minutes required to generate music for a one-minute video. All other metrics represent average rankings, where 1 is the best possible and 3 is the worst possible ranking. Across all metrics, lower is better.
4 Experimental setup
As discussed in Section 2.3.1, conducting an objective evaluation is challenging when using unmatched video and MIDI datasets. We therefore focus on the subjective evaluation of our overall model. We conduct a user study to compare our results with two open-source state-of-the-art models that generate MIDI from arbitrary videos: Video2Music Kang et al. (2024) and CMT Di et al. (2021).
4.1 Dataset
We train our music generator on the Lakh Pianoroll-5 dataset Dong et al. (2018) with valence-arousal labels Sulun et al. (2022). For evaluation, we perform inference on videos from the Pittsburgh Advertisements Dataset (Ads) Hussain et al. (2017). We select this dataset because advertisements often rely on music to maximize viewer engagement Zander (2006). Additionally, since none of the compared methods have used this dataset, it ensures a fair comparison. Finally, the Ads dataset includes sentiment labels, allowing us to construct a well-balanced test set in terms of emotion.
We first filter the dataset by selecting videos associated with the four basic emotions commonly used in music emotion classification: cheerful, calm, angry, and sad Yang et al. (2008). We then remove videos shorter than 1 minute. To ensure an unbiased selection, we use YouTube IDs, which are generated randomly by YouTube. We sort these IDs alphabetically and select the first six videos from each emotion category, resulting in a total of 24 test videos. Finally, we trim all videos to a uniform duration of 1 minute.
4.2 Implementation details
We use the Music Transformer model with relative global attention in our music generator Huang et al. (2018). The model consists of 11 layers, 8 attention heads, and a dimensionality of 512. We train it with a context length of 1216 and a batch size of 64. Training is performed using cross-entropy loss with a learning rate of 2e-4 for the first 300k steps, followed by 5e-5 for the next 300k steps. We use the Adam optimizer with gradient clipping to a norm of 1 Kingma and Ba (2015). We do not apply any regularization methods and did not observe overfitting, as the model is trained on large-scale, densely labeled data where it predicts each token of its input sequence Sulun (2018).
For data augmentation, we transpose the pitches of all instruments, except drums, by a randomly chosen integer between -3 and 3, inclusive. We set our model size to be comparable to the baseline methods, with our model containing 37M parameters, compared to 39M in CMT Di et al. (2021) and 33M in Video2Music Kang et al. (2024). We set the maximum absolute value of means across all emotion categories to 0.8. For all models, we synthesize MIDI files into audio waveforms using Fluidsynth [^2] https://www.fluidsynth.com and apply peak audio normalization up to -3 dB.
4.3 Subjective evaluation
Using the 24 test videos, we generate accompanying music with the three models and create 12 survey pages, each containing two videos. For each video, the three music versions, generated by the compared models, are presented side by side with anonymized model names, with each model’s output appearing equally in the left, center, and right positions. We enroll 36 participants: 12 with self-reported knowledge of music theory, forming Group 1, and 24 without, forming Group 2. Each survey page and each video are viewed by one participant from Group 1 and two participants from Group 2.
We ask participants to rank the three models based on the standard criteria used in previous works Di et al. (2021); Kang et al. (2024): Music richness (MR): The richness and diversity of the music, independent of the video content. Music quality (MQ): The overall quality of the music, independent of the video content. Emotion match (EM): How well the music matches the video in terms of emotion. Timing match (TM): How well the music synchronizes with the video in terms of rhythm and timing. Overall match (OM): How well the music matches the video as a whole. Each model is assigned a unique ranking of 1 (best), 2, or 3 (worst), ensuring no ties. Finally, we report runtimes (RT) as the average time (in minutes) required to generate music for a one-minute video, including model initialization.
5 Results and conclusion
Table 2 compares EMSYNC, Video2Music Kang et al. (2024), and CMT Di et al. (2021) across all subjective criteria based on average rankings. A lower score represents a better ranking, with 1 being the best and 3 the worst possible.
EMSYNC consistently outperforms both baselines across all metrics for both Group 1 (music theory-aware) and Group 2. Group 2 rated EMSYNC even higher, particularly in music richness and emotion match, indicating that its strengths are even more apparent to general listeners. EMSYNC is also the fastest, with a runtime of 1.417 minutes, compared to 1.607 minutes for Video2Music and 3.554 minutes for CMT.
These results confirm that EMSYNC produces the most diverse and high-quality music with the best emotional alignment, rhythmic synchronization, and overall video compatibility while also being the most computationally efficient model.
By automatically generating soundtracks for user-provided videos, our work has the potential to streamline video production and enhance viewer engagement while offering a valuable framework for both machine learning researchers and multimedia content creators.

### v2

@@
 It follows a two-stage framework, where a pretrained video emotion classifier extracts emotional features, and a conditional music generator produces MIDI sequences guided by both emotional and temporal cues.
-We introduce boundary offsets, a novel temporal conditioning mechanism that enables the model to anticipate and align musical chords with scene cuts.
+We introduce boundary offsets, a novel temporal conditioning mechanism that enables the model to anticipate upcoming video scene cuts and align generated musical chords with them.
 Unlike existing models, our approach retains event-based encoding, ensuring fine-grained timing control and expressive musical nuances.
 We also propose a mapping scheme to bridge the video emotion classifier, which produces discrete emotion categories, with the emotion-conditioned MIDI generator, which operates on continuous-valued valence-arousal inputs.
-In subjective listening tests, EMSYNC outperforms state-of-the-art models across all subjective metrics, for music theory-aware participants as well as the general listeners.
+Our method outperforms state-of-the-art models in objective and subjective evaluations across different video datasets, demonstrating its effectiveness in generating music aligned to video both emotionally and temporally.
 1 Introduction
-The online distribution of user-generated multimedia content is expanding at an exponential rate thanks to affordable, high-quality recording equipment and video editing software Falkowski-Gilski and Uhl (2020).
-A key challenge in content creation is providing suitable soundtracks to enhance viewer engagement Buhler (2018).
+The online distribution of user-generated multimedia content is expanding at an exponential rate thanks to affordable, high-quality recording equipment and video editing software (Falkowski-Gilski and Uhl 2020).
+A key challenge in content creation is providing suitable soundtracks to enhance viewer engagement (Buhler 2018).
 However, unauthorized use of commercially published music infringes copyright, preventing monetization for creators on platforms like YouTube.
 Alternatives such as purchasing music, hiring composers, or searching for royalty-free tracks are often costly, time-consuming, or fail to ensure proper synchronization with video content.
-Automatic video-based music generation offers a promising solution to this problem Di et al.
-(2021); Kang et al.
-(2024); Zhuo et al.
-(2023).
-Existing approaches that generate music as audio waveforms lack editability Lam et al.
-(2024); Copet et al.
-(2024).
-This method compresses all stages of music production—composition, performance, editing, mixing, and mastering—into a single process, limiting creative control for professionals.
+Automatic video-based music generation offers a promising solution to this problem (Di et al.
+2021; Zhuo et al.
+2023; Kang, Poria, and Herremans 2024).
+Existing approaches that generate music as audio waveforms lack editability (Lam et al.
+2024; Copet et al.
+2024).
+These methods compress all stages of music production, including composition, performance, editing, mixing, and mastering, into a single process, which limits creative control for musicians and producers.
 By contrast, generating music in a symbolic format like MIDI (Musical Instrument Digital Interface) offers greater flexibility.
 MIDI functions as a digital music score, encoding instrument names, note pitches, durations, and velocities.
-Professionals can edit compositions, synthesize audio using digital audio workstations (DAWs), or record and refine performances.
+Musicians and producers can edit compositions, synthesize audio using digital audio workstations (DAWs), or record and refine performances.
 This work focuses on generating music in MIDI format from arbitrary videos using deep neural networks (DNNs).
@@
 A major challenge in training video-to-MIDI DNNs is the absence of large-scale paired video-MIDI datasets.
-Existing datasets are either domain-specific, such as pianist hand videos Koepke et al.
-(2020), or contain limited samples, around 1k Di et al.
-(2021); Zhuo et al.
-(2023).
-To address this, we develop a model that generates music for any video type, leveraging the Lakh MIDI Dataset—the largest available MIDI dataset with 176,581 samples—ensuring diverse and high-quality outputs Raffel (2016).
+Existing datasets are either domain-specific, such as pianist hand videos (Koepke et al.
+2020), or contain limited samples, around 1k (Di et al.
+2021; Zhuo et al.
+2023).
+To address this, we develop a model that generates music for any video type by leveraging the Lakh MIDI Dataset (Raffel 2016), the largest available collection with 176,581 samples, to ensure diverse and high-quality outputs.
 Since the Lakh MIDI dataset lacks corresponding videos, we adopt a two-stage approach: extracting video features relevant to music generation and using them as conditioning inputs.
+Drawing from musicology literature, we identify two essential aspects for effective video-based music generation: emotional alignment (Deutsch 2007) and temporal synchronization (Prendergast 1992).
 We name our model EMSYNC, as it aligns music with video by matching their emotions, and synchronizing their temporal boundaries.
-While our method is applicable to any selection of temporal boundary and emotion representation, we define musical boundaries using long-duration chords, video boundaries using scene cuts, and represent emotions with the valence-arousal model Russell (1980).
+While our method is applicable to any selection of temporal boundary and emotion representation, we define musical boundaries using long-duration chords, video boundaries using scene cuts, and represent emotions with the valence-arousal model (Russell 1980).
 Specifically, we extract scene cut locations from the input video and guide the music generator to produce long-duration chords near those locations that are rhythmically and harmonically compatible with the rest of the generated music.
-We also use a pretrained video emotion classifier to estimate discrete emotion probabilities Sulun et al.
-(2024a), map them to valence-arousal values, and condition our music generator accordingly.
+We also use a pretrained video emotion classifier to estimate discrete emotion probabilities, map them to valence-arousal values, and condition our music generator accordingly.
 Temporal conditioning in MIDI-generating models presents unique challenges.
-While deep transformer models can handle time-based data, such as videos, their sequence dimension correlates linearly with time due to a fixed frame rate Liu et al.
-(2022); Arnab et al.
-(2021).
-In contrast, MIDI is typically processed using an event-based representation, where sequence and time dimensions are not linearly correlated Oore et al.
-(2020).
+While deep transformer models can handle time-based data, such as videos, their sequence dimension correlates linearly with time due to a fixed frame rate (Arnab et al.
+2021; Liu et al.
+2022).
+In contrast, MIDI is typically processed using an event-based representation, where sequence and time dimensions are not linearly correlated (Oore et al.
+2020).
 In event-based MIDI encoding, two primary token types exist: “note” and “time shift”.
@@
 The key advantage of event-based encoding is the absence of a fixed time grid, allowing for subtle, expressive timing variations that reflect human musicianship.
-Unlike state-of-the-art video-based music generators that rely on fixed, coarse time grids Di et al.
-(2021); Kang et al.
-(2024), we introduce a novel temporal conditioning method that preserves event-based encoding, enabling fine-grained temporal control.
-We also tackle the challenge of using emotions as an intermediary to link music generation to video.
-Sulun et.
-al Sulun et al.
-(2022) previously assigned valence-arousal emotion labels to samples in the Lakh MIDI dataset Raffel (2016), creating a training dataset two orders of magnitude larger than existing emotion-labeled MIDI datasets Panda et al.
-(2013); Ferreira and Whitehead (2019); Hung et al.
-(2021).
-However, the largest emotion-labeled video datasets primarily provide categorical labels Xu et al.
-(2016); Sun et al.
-(2020).
-To bridge this gap, we employ the VEMOCLAP model (Video EMOtion Classifier using Pretrained features) Sulun et al.
-(2024a), trained on the Ekman-$6$ dataset Xu et al.
-(2016).
-Since VEMOCLAP outputs probabilities of categorical emotions, we map these to valence-arousal values using prior user study results on emotion Russell and Mehrabian (1977).
-We compare our method to state-of-the-art video-based MIDI generators through subjective listening tests, where it outperforms all competing methods across all metrics for both music theory-aware participants and general listeners while also being the fastest.
-Our key contributions include the following:
-  • We develop a video-based MIDI generator that surpasses state-of-the-art models across all subjective metrics and participant demographics.
+Unlike state-of-the-art video-based music generators that rely on fixed, coarse time grids (Di et al.
+2021; Kang, Poria, and Herremans 2024), we introduce a novel temporal conditioning method that preserves event-based encoding, enabling fine-grained temporal control.
+The selection of temporal features for video-music synchronization is also non-trivial.
+Existing work matches video motion speed and motion saliency with musical note density, producing a dense representation of temporal dynamics that leads to continuous conditioning and frequent changes in note density (Di et al.
+2021).
+Our inspection of the output samples from this approach identified two key limitations.
+First, the continuously varying note density can disrupt rhythmic consistency, making the underlying beat unstable and harder to follow.
+Second, because the mapping is dense, synchronization between music and video is difficult to perceive unless there are strong contrasts in motion speed and note density across consecutive sections.
+Motivated by works in musicology and soundtrack composition, we choose a temporal representation of video using sections, separated by sparse temporal boundaries of scene cuts (Cohen 1990; Prendergast 1992).
+To enable our music generator to anticipate future boundaries and align its output accordingly, we propose using boundary offsets.
+We conduct objective and subjective evaluations to measure the temporal synchronization between input video and generated music, and demonstrate that our approach outperforms dense temporal conditioning.
+In addition, we tackle the challenge of using emotions as an intermediary to link music generation to video.
+Emotions are commonly represented with either categorical (Ekman 1971) or dimensional (continuous-valued) (Russell 1980) models.
+Based on prior findings from psychology and emotion research (Russell and Mehrabian 1977), we provide a mapping between categorical and dimensional emotion representations.
+This mapping enables the combined use of different emotion representations across data modalities, such as categorical emotions for video classification and dimensional emotions for MIDI generation, as in our case.
+This approach ensures flexibility and supports the use of training data labeled with either type of emotion representation.
+We compare our method to state-of-the-art video-based MIDI generators through objective and subjective evaluation on different datasets, where it outperforms the state-of-the-art methods.
+Our main contributions are as follows:
+  • We develop a video-based MIDI generator that surpasses state-of-the-art models across all subjective and a majority of the objective metrics for all datasets.
   • We introduce boundary offsets for temporal conditioning in transformers, ensuring precise alignment between the input video and the output MIDI.
-  • We leverage insights from psychological studies to map discrete emotion categories to continuous valence-arousal values, enabling an emotional connection between the input video and the generated music.
-  • We design our model architecture to modify only the inputs to the transformer body, allowing future users to fine-tune it for other tasks and data structures seamlessly.
+  • Based on psychological research, we establish a mapping between discrete emotion categories and continuous valence-arousal values, an enabling emotional connection between the input video and the generated music.
 2 Related work
-In this section, we introduce the video emotion classifier and the emotion-based MIDI generator used in our work.
-We then review existing video-based MIDI generation methods and discuss their evaluation approaches.
-2.1 Video emotion classification
-The VEMOCLAP model exploits publicly available pretrained models for a multimodal video analysis and uses the resulting pretrained features to classify the emotions of arbitrary videos Sulun et al.
-(2024a, b).
-It employs pretrained models for automatic speech recognition (ASR), optical character recognition (OCR), facial expression classification, audio classification, and image understanding.
-Textual features obtained from ASR and OCR are further processed using a text sentiment classifier.
-Cross-attention layers Vaswani et al.
-(2017) integrate and process these multimodal features, and a linear layer produces the final emotion probabilities.
-The model is trained on the Ekman-6 dataset Xu et al.
-(2016), consisting of 1637 videos labeled with the six basic emotions derived from the original work of Ekman: anger, disgust, fear, joy, sadness, and surprise Ekman (1971).
-2.2 Emotion-based MIDI generation
-Sulun et al.
-Sulun et al.
-(2022) collected valence and arousal values for songs in the Lakh Pianoroll Dataset Dong et al.
-(2018) and trained a conditional transformer using the resulting labeled dataset.
-To integrate continuous-valued valence and arousal features with discrete musical note tokens, they project valence and arousal into a vector space using separate linear layers.
-Musical input tokens are projected into vectors of the same dimensionality using an embedding layer.
-These vectors are then concatenated with the projected valence and arousal vectors along the sequence dimension and fed into the transformer body with relative global attention Huang et al.
-(2018).
-2.3 Video-based MIDI generation
 While our method applies to arbitrary videos, several works focus on generating symbolic music for specific types of videos, such as those featuring human movements like dancing or instrumental performances.
-The Foley Music model Gan et al.
-(2020) generates MIDI from videos of musicians by processing body keypoint movements using a Graph Convolutional Network Kipf and Welling (2017) and a Transformer Vaswani et al.
-(2017).
+The Foley Music model (Gan et al.
+2020) generates MIDI from videos of musicians by processing body keypoint movements using a Graph Convolutional Network (Kipf and Welling 2017) and a Transformer (Vaswani et al.
+2017).
 Similarly, Koepke et al.
-Koepke et al.
 (2020) and Su et al.
-Su et al.
-(2020) utilize use deep neural networks with residual connections to generate symbolic music from videos of finger movements on piano keyboards.
-Due to their specialized nature, these approaches rely on datasets containing video-MIDI pairs, though these datasets typically contain fewer than 1k samples Zhao et al.
-(2018); Li et al.
-(2018); Koepke et al.
-(2020); Su et al.
-(2020).
-The RhythmicNet model employs a multi-stage process to generate music from dance videos by predicting beats and style, generating a drum track, and subsequently creating multitrack music Su et al.
-(2021).
+(2020) use deep neural networks with residual connections to generate symbolic music from videos of finger movements on piano keyboards.
+Due to their specialized nature, these approaches rely on datasets containing video-MIDI pairs, however, these datasets typically contain fewer than 1k samples (Zhao et al.
+2018; Li et al.
+2018; Koepke et al.
+2020; Su, Liu, and Shlizerman 2020).
+The RhythmicNet model employs a multi-stage process to generate music from dance videos by predicting beats and style, generating a drum track, and subsequently creating multitrack music (Su, Liu, and Shlizerman 2021).
 Some studies explore the more general task of generating symbolic music for arbitrary videos.
-The most similar to our approach, the Controllable Music Transformer (CMT), generates music based on video features such as motion speed, motion saliency, and timing Di et al.
-(2021).
-CMT employs the Lakh Pianoroll Dataset Dong et al.
-(2018), the same dataset we use, and processes music using an extended Compound Word representation, where each token encodes type, beat/bar marking, note strength, note density, instrument, pitch, and duration Hsiao et al.
-(2021).
+The most similar to our approach, the Controllable Music Transformer (CMT), generates music based on video features such as motion speed, motion saliency, and timing (Di et al.
+2021).
+CMT employs the Lakh Pianoroll Dataset (Dong et al.
+2018), the same dataset we use, and processes music using an extended Compound Word representation, where each token encodes type, beat/bar marking, note strength, note density, instrument, pitch, and duration (Hsiao et al.
+2021).
 While this representation reduces sequence length, it significantly increases input dimensionality.
 Consequently, CMT uses a low temporal resolution of four ticks per beat, translating to a time resolution of 125 milliseconds for a song at 120 beats per minute.
-Furthermore, note events align precisely with beat subdivisions, whereas human musicians introduce expressive timing deviations.
-In contrast, our model utilizes an event-based representation with independent time shift tokens at an 8-millisecond resolution.
-This resolution is sufficiently fine to capture musical nuance while maintaining absolute time rather than rigid beat subdivisions Oore et al.
-(2020).
+Furthermore, note events align precisely with beat subdivisions, whereas human performance naturally deviates from beat subdivisions due to expressive timing (Fabian, Timmers, and Schubert 2014).
+In contrast, our model utilizes an event-based representation with independent time shift tokens at an 8 ms resolution.
+This resolution is sufficient to capture musical nuance while maintaining absolute time rather than rigid beat subdivisions (Oore et al.
+2020).
 Additionally, unlike CMT, our model integrates high-level emotional conditioning alongside low-level temporal features.
 In a follow-up to CMT, Zhuo et al.
-introduced the Symbolic Music Videos dataset, containing video-MIDI pairs Zhuo et al.
-(2023).
-They sourced MIDI data from piano tutorial videos, automatically transcribing audio using the Onsets and Frames model, resulting in 1140 samples Hawthorne et al.
-(2018).
-Video features such as color histograms, RGB frame differences for motion, and high-level CLIP features Radford et al.
-(2021) were extracted to train three models sequentially: one generating chord sequences, another generating the melody, and a third generating the accompaniment.
-The Video2Music model shares similarities with our approach by utilizing both low-level video features and high-level emotional conditioning but differs in application Kang et al.
-(2024).
+introduced the Symbolic Music Videos dataset, containing video-MIDI pairs (2023).
+They sourced MIDI data from piano tutorial videos, automatically transcribing audio using the Onsets and Frames model, resulting in 1140 samples (Hawthorne et al.
+2018).
+Video features such as color histograms, RGB frame differences for motion, and high-level CLIP features (Radford et al.
+2021) were extracted to train three models sequentially: one generating chord sequences, another generating the melody, and a third generating the accompaniment.
+The Video2Music model shares similarities with our approach by utilizing both low-level video features and high-level emotional conditioning, but differs in application (Kang, Poria, and Herremans 2024).
 The authors compiled the MuVi-Sync dataset, consisting of 748 music videos labeled with musical attributes like note density, loudness, chords, and key.
@@
 Additionally, its use of fixed arpeggiation patterns limits musical diversity, and requiring user-defined chord and key inputs restricts accessibility for non-musicians.
-2.3.1 Evaluation methods
-Evaluation of video-based music generators is non-trivial.
-For objective evaluation, most works measure the difference between generated and ground-truth samples using metrics such as Pitch Class Histogram Entropy, Grooving Pattern Similarity, and Structureness Indicator Wu and Yang (2020).
-When the dataset contains paired video and MIDI samples, this method is straightforward Gan et al.
-(2020); Zhuo et al.
-(2023).
-For unpaired video and MIDI samples, objectively evaluating a model that generates MIDI from arbitrary videos is not feasible.
-The authors of the Controllable Music Transformer, using unpaired datasets, evaluate output MIDIs generated unconditionally, without input video Di et al.
-(2021).
-They empirically show that unconditionally generated MIDI samples exhibit metrics closer to the unpaired MIDI dataset compared to video-conditioned output samples.
-This occurs because constraining the model with video structure causes deviations from intrinsic MIDI structures, and the best way to match a specific MIDI dataset’s metrics is to train exclusively on that dataset.
-Subjective evaluation through user studies remains the most common, and often the only, viable method for assessing video-based MIDI generators Di et al.
-(2021); Kang et al.
-(2024); Zhuo et al.
-(2023); Gan et al.
-(2020); Su et al.
-(2021).
 3 Methodology
@@
 3.1 Conditional music generator
-Our music generator is conditioned on emotions as valence-arousal values and temporal boundaries.
-Figure 2 illustrates the model, with the emotion-conditioning mechanism in the upper part and the boundary-conditioning mechanism in the lower part.
+We extend an existing emotion-based MIDI generator trained with valence-arousal labels (Sulun, Davies, and Viana 2022) assigned to the Lakh MIDI dataset (Raffel 2016).
+By introducing temporal boundary conditioning and emotion mapping, we align music generation with input videos at inference time.
+Figure 2 illustrates the music generator, with the emotion-conditioning mechanism shown in the upper part and the boundary-conditioning mechanism in the lower part.
 Figure 1: Video-based music generation pipeline. The text and the image next to the arrows demonstrate sample values.
-3.1.1 Training dataset and preprocessing
-We train our music generator on the Lakh Pianoroll Dataset (LPD) Dong et al.
-(2018), which contains 174,154 pianorolls derived from the Lakh MIDI Dataset Raffel (2016).
-We tokenize the pianorolls using an event-based symbolic music representation Oore et al.
-(2020).
+Training dataset and preprocessing
+We train our music generator on the Lakh Pianoroll Dataset (LPD) (Dong et al.
+2018), which contains 174,154 pianorolls derived from the Lakh MIDI Dataset (Raffel 2016).
+We tokenize the pianorolls using an event-based symbolic music representation (Oore et al.
+2020).
 Specifically, an ON (note on) token marks the start of a note, and an OFF (note off) token marks its end.
@@
 For example, a piano note with a MIDI pitch of 60 (C4) is denoted as PIANO_ON_60.
-Since a larger number of instruments increases vocabulary size, we use the Lakh Pianoroll Dataset-5 variant, where all instrument tracks are merged into five predefined categories: bass, drums, guitar, piano, and strings Dong et al.
-(2018).
+Since a larger number of instruments increases vocabulary size, we use the Lakh Pianoroll Dataset-5 variant, where all instrument tracks are merged into five predefined categories: bass, drums, guitar, piano, and strings (Dong et al.
+2018).
 However, our method is adaptable to datasets with different instrument groupings.
@@
 Each token specifies a time increment in milliseconds.
-For example, an 800-millisecond shift is encoded as TIMESHIFT_800.
-We use a temporal resolution of 8 milliseconds with a maximum shift of 1000 milliseconds.
+For example, an 800 ms shift is encoded as TIMESHIFT_800.
+We use a temporal resolution of 8 ms with a maximum shift of 1000 ms.
 Longer durations are represented using multiple consecutive TIMESHIFT tokens.
 We also use the START tokens to mark the beginning of the songs, the BAR tokens to indicate the musical bars, and the PAD tokens to standardize input sequence lengths in minibatches.
-We use the CHORD token to mark long-duration chords as temporal boundaries, as they often serve as anchor points in musical compositions Bharucha and Krumhansl (1983).
+We introduce the CHORD token to mark long-duration chords as temporal boundaries, as they often serve as anchor points in musical compositions (Bharucha and Krumhansl 1983).
 However, this is a design choice, and our method can accommodate any selection of temporal boundaries.
-For emotion-conditioned generation, we use the model architecture and valence-arousal labeled dataset of Sulun et al., detailed in Section 2.2, which includes 23,661 samples Sulun et al.
-(2022).
-To utilize the remaining unlabeled samples, we further modify the model to accept NaN (Not A Number) values for valence or arousal, enabling both conditional and unconditional generation.
+To fully exploit available training data, we implement several strategies.
+Since we aim to generate multi-instrument compositions, we prioritize pieces with three or more instruments.
+However, one third of the Lakh MIDI dataset contains songs with only one or two instruments, with classical piano pieces dominating the single-instrument subset (Raffel 2016).
+As these samples still contain valuable information on musical structure, we avoid filtering them out.
+Instead, we prepend special tokens: FEWER_INSTRUMENTS for songs with two or fewer instruments and MORE_INSTRUMENTS for those with three or more.
+These tokens allow users to specify instrumentation preferences at inference time while leveraging the entire dataset during training.
+Additionally, to utilize MIDI samples without valence-arousal labels, we allow the model to accept NaN (Not A Number) values for valence or arousal, enabling both conditional and unconditional generation.
 If valence or arousal is unspecified, we assign it a NaN value and use a learned vector in place of its projection vector.
 Figure 2: Our music generator. Numbers underneath valence and arousal are sample values.
-3.1.2 Temporal boundary matching
-In addition to matching emotions, we also temporally synchronize the input video and output music by aligning their temporal boundaries.
-We define temporal boundaries as scene cut locations in video and chord locations in music.
+Temporal boundary matching
+Our model generates music that is synchronized with the video’s temporal boundaries.
+We define temporal boundaries as scene cut locations in video and long-duration chord locations in music.
 Since we employ a hybrid approach, we first train a music generator capable of incorporating boundary locations as input and producing music with chords near these boundaries.
@@
 In training, the temporal locations of chords from the ground-truth MIDI serve as input boundaries, whereas during video-based inference, we replace these boundaries with video scene cut locations.
-Figure 3: Graphical illustration of boundaries.
-Top: In symbolic music, a chord with three or more simultaneous notes and a duration exceeding a set threshold defines a musical boundary and is used during training.
-Middle: In video, scene cuts serve as video boundaries and are used during inference.
-Bottom: Boundary offset represents the temporal distance to the next boundary.
-These figures are illustrative; offsets do not perfectly align with the music except at the boundary.
-Chords are integral to a melody, providing harmonic and rhythmic support to surrounding notes, both preceding and following Bharucha and Krumhansl (1983).
+Chords are integral to a melody, providing harmonic and rhythmic support to surrounding notes, both preceding and following (Bharucha and Krumhansl 1983).
 During inference, forcing a CHORD token into the sequence at a specific location may cause the chord to sound off-beat or overly abrupt.
@@
 Sample offsets are illustrated in the bottom row of Figure 3.
-We process boundary offsets through a feed-forward network, generating boundary offset encodings, which are then concatenated with learned positional encodings Wang et al.
-(2020) along the feature dimension, as shown in the lower part of Figure 2.
+We process boundary offsets through a feed-forward network, generating boundary offset encodings, which are then concatenated with learned positional encodings (Wang et al.
+2020) along the feature dimension, as shown in the lower part of Figure 2.
 This architectural choice is motivated by several key factors.
@@
 Second, we avoid adding boundary offset encodings directly to learned positional encodings to maintain a distinction between the two.
-Finally, recent studies suggest that decoder-only transformers can implicitly learn positional encodings through their internal weights, even without explicitly adding them Haviv et al.
-(2022); Kazemnejad et al.
-(2024).
+Finally, recent studies suggest that decoder-only transformers can implicitly learn positional encodings through their internal weights, even without explicitly adding them (Haviv et al.
+2022; Kazemnejad et al.
+2024).
 Based on this insight, we halve the feature length of learned positional encodings and allocate the remaining feature space to boundary offset encodings.
 The resulting vector sequence consists of positional encodings augmented with boundary offset encodings.
-Following the standard transformer model, we add this sequence to the token embeddings Vaswani et al.
-(2017) before passing it to the transformer body with relative global attention Huang et al.
-(2018).
+Following the standard transformer model, we add this sequence to the token embeddings (Vaswani et al.
+2017) before passing it to the transformer body with relative global attention (Huang et al.
+2018).
 The transformer’s input is represented as:
@@
 The operations $\text{Concat}_{s}$ and $\text{Concat}_{f}$ denote concatenation along the sequence and feature dimensions, respectively.
-Our music generator architecture is illustrated in Figure 2.
+Figure 3: Graphical illustration of boundaries.
+Top: In symbolic music, a chord with three or more simultaneous notes and a duration exceeding a set threshold defines a musical boundary and is used during training.
+Middle: In video, scene cuts serve as video boundaries and are used during inference.
+Bottom: Boundary offset represents the temporal distance to the next boundary.
+These figures are illustrative; offsets do not perfectly align with the music except at the boundary.
 Algorithm 1 outlines the process of computing boundary offsets during inference, i.e., music generation.
 The model generates music autoregressively, producing one token per forward pass.
-By tracking the generated TIMESHIFT tokens, we maintain a time cursor and compute the boundary offset, i.e., the time remaining until the next boundary, for each generated token.
+We maintain a time cursor by tracking the generated TIMESHIFT tokens and compute the boundary offset, i.e., the time remaining until the next boundary, for each generated token.
 If the model generates a CHORD token, we calculate the absolute difference between the time cursor and each input boundary.
-The boundary with a difference smaller than the predefined sensitivity threshold of 1 second is considered successfully generated.
+The boundary with a difference smaller than the predefined sensitivity threshold of 1 s is considered successfully generated.
 We then remove these boundaries from future offset calculations by replacing them with infinity.
@@
   end if
-     $\boldsymbol{\delta}$.append$(\max(\min(\mathbf{b}-c),\delta_{max}))$
-     $\mathbf{t}$.append$((t_{t},t_{v}))$
+[…]
+[…]
 end while
 3.2 Scene detection
-During video-based inference, we first extract the video’s temporal boundaries, specifically its scene cut locations, using the FFmpeg software [^1] https://www.ffmpeg.org/.
-We then apply a difference filter to the extracted scene cuts and remove those occurring less than 4 seconds apart.
+During video-based inference, we first extract the video’s temporal boundaries, specifically its scene cut locations.
+As scene detection is a solved problem, we employ the off-the-shelf FFmpeg software [^1] https://www.ffmpeg.org/.
+As we generate musical chords near scene boundaries, we aim to avoid an excessive number of densely spaced chords.
+We therefore apply a difference filter to the extracted scene cuts and remove those occurring less than 4 s apart.
 The remaining timestamps serve as the input boundaries for the music generator.
-3.3 Emotion matching
-We use the VEMOCLAP emotion classifier to extract emotions from input videos as probability distributions over discrete emotion categories Sulun et al.
-(2024a).
-The music generator is conditioned on emotions represented as valence and arousal values ranging from -1 to 1, rather than discrete categories Sulun et al.
-(2022).
-Since the VEMOCLAP emotion classifier outputs discrete emotion categories, we first map them to the valence-arousal plane Russell (1980) before feeding them into our music generator.
-Russell and Mehrabian conducted user studies to find the corresponding valence and arousal values of discrete emotion categories Russell and Mehrabian (1977).
+3.3 Emotion mapping
+We use a pretrained emotion classifier to extract emotions from input videos as probability distributions over discrete emotion categories (Sulun, Viana, and Davies 2024).
+Our music generator is conditioned on emotions represented as valence and arousal values ranging from -1 to 1, rather than discrete categories.
+Since the emotion classifier outputs discrete emotion categories, we first map them to the valence-arousal plane (Russell 1980) before feeding them into the music generator.
+Russell and Mehrabian conducted user studies to find the corresponding valence and arousal values of discrete emotion categories (1977).
 They presented their findings in a table that contains the mean and standard deviation values of valence and arousal for each categorical emotion.
-We extracted and used the values corresponding to Ekman’s six basic emotions, which are presented in Table 1 Ekman (1971).
+We extracted and used the values corresponding to Ekman’s six basic emotions (1971), which are presented in Table 1.
 Valence Arousal
@@
 surprise -0.40 0.30 -0.67 0.27
-Table 1: Means and standard deviations of valence and arousal values corresponding to Ekman’s categorical emotions Ekman (1971), obtained from the user studies of Russell and Mehrabian Russell and Mehrabian (1977).
+Table 1: Means and standard deviations of valence and arousal values corresponding to Ekman’s categorical emotions (1971), obtained from the user studies of Russell and Mehrabian (1977).
 Using the output probabilities of the video classifier along with the means and standard deviations of each emotion, we construct a mixture of Gaussian distributions.
-Users can either sample from the mixture—first selecting an emotion category based on the output probabilities and then sampling from the corresponding Gaussian distribution—or use the mean of the mixture, which is computed as the weighted average of the emotion category means, with weights determined by the output probabilities.
+Valence-arousal values can be obtained either by sampling an emotion category based on the output probabilities and then drawing from its corresponding Gaussian, or by using the mixture mean computed as the weighted average of the category means.
 Additionally, we introduce a parameter that represents the maximum absolute value of means across all emotion categories.
@@
 This approach allows users to adjust the conditioning of the music generator, choosing between a wider or narrower range of emotions.
-Group 1 Group 2
-MR MQ EM TM OM MR MQ EM TM OM RT
-Video2Music 2.521 1.854 2.125 1.958 2.167 2.708 1.771 1.854 1.979 2.083 1.607
-CMT 1.854 2.438 2.000 2.125 2.083 1.792 2.562 2.458 2.312 2.354 3.554
-EMSYNC (Ours) 1.625 1.708 1.875 1.917 1.750 1.500 1.667 1.688 1.708 1.562 1.417
-Table 2: Group 1 consists of participants with a self-reported understanding of music theory, while Group 2 includes those without it.
-We use the following abbreviations: MR (music richness), MQ (music quality), EM (emotion match), TM (timing match), OM (overall match), and RT (runtime).
-Runtime measures the minutes required to generate music for a one-minute video.
-All other metrics represent average rankings, where 1 is the best possible and 3 is the worst possible ranking.
-Across all metrics, lower is better.
+Although not used in our work, it is possible to invert this mapping, that is, to map valence-arousal values to emotion categories by comparing the valence-arousal coordinates to each category’s distribution defined by its mean and standard deviation.
+This involves calculating a distance metric, such as Euclidean or Mahalanobis distance, or computing the likelihood assuming Gaussian distributions, and then assigning the category with the smallest distance or highest likelihood.
 4 Experimental setup
-As discussed in Section 2.3.1, conducting an objective evaluation is challenging when using unmatched video and MIDI datasets.
-We therefore focus on the subjective evaluation of our overall model.
-We conduct a user study to compare our results with two open-source state-of-the-art models that generate MIDI from arbitrary videos: Video2Music Kang et al.
-(2024) and CMT Di et al.
-(2021).
-4.1 Dataset
-We train our music generator on the Lakh Pianoroll-5 dataset Dong et al.
-(2018) with valence-arousal labels Sulun et al.
-(2022).
-For evaluation, we perform inference on videos from the Pittsburgh Advertisements Dataset (Ads) Hussain et al.
-(2017).
-We select this dataset because advertisements often rely on music to maximize viewer engagement Zander (2006).
-Additionally, since none of the compared methods have used this dataset, it ensures a fair comparison.
-Finally, the Ads dataset includes sentiment labels, allowing us to construct a well-balanced test set in terms of emotion.
-We first filter the dataset by selecting videos associated with the four basic emotions commonly used in music emotion classification: cheerful, calm, angry, and sad Yang et al.
-(2008).
+We perform objective and subjective evaluations of our overall model using two datasets, against the open-source state-of-the-art models that generate MIDI from arbitrary videos: Video2Music (Kang, Poria, and Herremans 2024) and CMT (Di et al.
+2021).
+4.1 Implementation details
+Our music generator consists of 11 layers, 8 attention heads, and a model dimensionality of 512.
+To ensure a fair comparison, we align our model size with the compared models: our model has 37M parameters, compared to 39M in CMT (Di et al.
+2021) and 33M in Video2Music (Kang, Poria, and Herremans 2024).
+Training is done with a context length of 1216 and a batch size of 64.
+The model is implemented in PyTorch (Paszke et al.
+2019) and trained on a single NVIDIA A100 80GB GPU.
+For video-based inference in evaluation, we set the maximum absolute value of the emotion category means to 0.8.
+To ensure consistency, all methods generate MIDI outputs that are synthesized into audio waveforms using Fluidsynth [^2] https://www.fluidsynth.com, followed by peak normalization to -3 dB.
+The audio is then merged with the input video by overwriting its audio channel.
+For all methods, outputs are generated in a single run across all datasets using a fixed random seed.
+We train our model using cross-entropy loss with a learning rate of 2e-4 for the first 300k steps, followed by 5e-5 for the next 300k steps.
+We use the Adam optimizer with gradient clipping to a norm of 1 (Kingma and Ba 2015).
+For data augmentation, we transpose the pitches of all instruments, except drums, by a randomly chosen number of semitones between -3 and 3, inclusive.
+We reserve 5% of the training data for validation.
+Hyperparameters are selected via grid search to minimize validation loss.
+4.2 Evaluation datasets
+For objective and subjective evaluation, we first use the EmoMV-C dataset, which consists of music videos containing music audio tracks to be used as ground truth for objective evaluation (Thao, Roig, and Herremans 2023).
+We exclude auditory features during the emotion classification to prevent data leakage.
+To assess the model’s generalization across different video types, we also include the Pittsburgh Advertisements (Ads) dataset, which contains advertisement videos that may or may not feature music (Hussain et al.
+2017).
+As advertisements often rely on music to maximize viewer engagement, the Ads dataset provides a real-world scenario for evaluation.
+(Zander 2006).
+We use the full EmoMV-C validation split, which contains 48 thirty-second videos.
+We filter the Ads dataset to match the total duration of EmoMV-C.
+We select videos associated with the four basic emotions commonly used in music emotion classification: cheerful, calm, angry, and sad (Yang et al.
+2008).
 We then remove videos shorter than 1 minute.
 To ensure an unbiased selection, we use YouTube IDs, which are generated randomly by YouTube.
-We sort these IDs alphabetically and select the first six videos from each emotion category, resulting in a total of 24 test videos.
-Finally, we trim all videos to a uniform duration of 1 minute.
-4.2 Implementation details
-We use the Music Transformer model with relative global attention in our music generator Huang et al.
-(2018).
-The model consists of 11 layers, 8 attention heads, and a dimensionality of 512.
-We train it with a context length of 1216 and a batch size of 64.
-Training is performed using cross-entropy loss with a learning rate of 2e-4 for the first 300k steps, followed by 5e-5 for the next 300k steps.
-We use the Adam optimizer with gradient clipping to a norm of 1 Kingma and Ba (2015).
-We do not apply any regularization methods and did not observe overfitting, as the model is trained on large-scale, densely labeled data where it predicts each token of its input sequence Sulun (2018).
-For data augmentation, we transpose the pitches of all instruments, except drums, by a randomly chosen integer between -3 and 3, inclusive.
-We set our model size to be comparable to the baseline methods, with our model containing 37M parameters, compared to 39M in CMT Di et al.
-(2021) and 33M in Video2Music Kang et al.
-(2024).
-We set the maximum absolute value of means across all emotion categories to 0.8.
-For all models, we synthesize MIDI files into audio waveforms using Fluidsynth [^2] https://www.fluidsynth.com and apply peak audio normalization up to -3 dB.
+We sort these IDs alphabetically and select the first six videos from each emotion category, resulting in 24 evaluation videos.
+Finally, we trim the videos to a uniform duration of 1 minute.
+The resulting evaluation videos are then fed into the compared models.
+4.3 Objective evaluation
+We first evaluate music similarity using the ground-truth audio from the EmoMV-C dataset.
+Although the ground truth is not in MIDI, all methods generate MIDI outputs that are synthesized to audio, enabling a fair comparison in the audio domain.
+We evaluate the Kullback-Leibler Divergence (KL) between the labels of generated and ground-truth music, calculated using a pretrained music tagging model (Won et al.
+2020).
+This measures how closely the distribution of semantic tags in the generated music matches that of the ground-truth.
+We avoid using Fréchet Audio Distance, as it compares overall audio distributions suited for unconditional generation (Kilgour et al.
+2019), whereas video-based music generation requires per-sample evaluation to assess alignment with each video’s unique content.
+We additionally evaluate temporal audio-video alignment (AV-Align) score, which is the Intersection-over-Union between video and audio peaks (Yariv et al.
+2024), using both EmoMV-C and Ads datasets.
+Video peaks are detected from the mean optical flow magnitude (Horn and Schunck 1981), while audio peaks correspond to audio onsets (Böck and Widmer 2013).
+We report AV-Align scores using windows of one video frame (33.4 ms at 30 FPS) and one second.
+Each window is applied to detect local peaks in both video and audio and then used to match peaks within the same temporal range for alignment.
 4.4 Subjective evaluation
-Using the 24 test videos, we generate accompanying music with the three models and create 12 survey pages, each containing two videos.
-For each video, the three music versions, generated by the compared models, are presented side by side with anonymized model names, with each model’s output appearing equally in the left, center, and right positions.
-We enroll 36 participants: 12 with self-reported knowledge of music theory, forming Group 1, and 24 without, forming Group 2.
-Each survey page and each video are viewed by one participant from Group 1 and two participants from Group 2.
-We ask participants to rank the three models based on the standard criteria used in previous works Di et al.
-(2021); Kang et al.
-(2024): Music richness (MR): The richness and diversity of the music, independent of the video content.
-Music quality (MQ): The overall quality of the music, independent of the video content.
-Emotion match (EM): How well the music matches the video in terms of emotion.
-Timing match (TM): How well the music synchronizes with the video in terms of rhythm and timing.
-Overall match (OM): How well the music matches the video as a whole.
-Each model is assigned a unique ranking of 1 (best), 2, or 3 (worst), ensuring no ties.
-Finally, we report runtimes (RT) as the average time (in minutes) required to generate music for a one-minute video, including model initialization.
+Using the evaluation videos from the EmoMV-C and the Ads datasets, we generate accompanying music with the three models and create survey pages, each containing two videos.
+For each video, the three music versions generated by the compared models are presented side by side with anonymized model names, with each model’s output appearing equally in the left, center, and right positions.
+We also provide output samples from eight videos online, with model names deanonymized and each model’s output appearing in the same position for clarity.
+[^3] https://em-sync.github.io
+We enrolled 153 participants and asked them to rank the three models based on the standard criteria used in previous works (Di et al.
+2021; Kang, Poria, and Herremans 2024): Music Richness (MR) — the richness and diversity of the music, independent of the video content; Music Quality (MQ) — the overall quality of the music, independent of the video content; Emotion Match (EM) — how well the music matches the video in terms of emotion; Timing Match (TM) — how well the music synchronizes with the video in terms of rhythm and timing; and Overall Match (OM) — how well the music matches the video as a whole.
+Each model is assigned a unique ranking of 1 (best), 2, or 3 (worst).
 5 Results and conclusion
-Table 2 compares EMSYNC, Video2Music Kang et al.
-(2024), and CMT Di et al.
-(2021) across all subjective criteria based on average rankings.
-A lower score represents a better ranking, with 1 being the best and 3 the worst possible.
-EMSYNC consistently outperforms both baselines across all metrics for both Group 1 (music theory-aware) and Group 2.
-Group 2 rated EMSYNC even higher, particularly in music richness and emotion match, indicating that its strengths are even more apparent to general listeners.
-EMSYNC is also the fastest, with a runtime of 1.417 minutes, compared to 1.607 minutes for Video2Music and 3.554 minutes for CMT.
-These results confirm that EMSYNC produces the most diverse and high-quality music with the best emotional alignment, rhythmic synchronization, and overall video compatibility while also being the most computationally efficient model.
+Table 2 presents objective evaluation results for EMSYNC, Video2Music (Kang, Poria, and Herremans 2024), and CMT (Di et al.
+2021), on the EmoMV-C and Ads datasets.
+EMSYNC outperforms both baselines across a majority of the metrics.
+On EmoMV-C, it achieves the best frame-level alignment and the lowest KL divergence, indicating superior synchronization to video and similarity to ground-truth music.
+While Video2Music performs best on second-level alignment, EMSYNC remains competitive.
+On the Ads dataset, EMSYNC obtains the best scores across both alignment metrics, further demonstrating its robustness across different video domains.
+We don’t calculate KL divergence because the Ads dataset doesn’t include ground truth music.
+Table 3 demonstrates the subjective results in terms of average user rankings with 95% confidence margins.
+A lower average score represents a better ranking, with 1 being the best and 3 the worst possible.
+EMSYNC achieves the best rankings across all metrics and datasets.
+The difference is more pronounced in video-music correspondence metrics such as emotion match and timing match, highlighting our model’s stronger alignment to video.
+For overall match, our method outperforms the others with non-overlapping 95% confidence intervals, reflecting a decisive user preference for EMSYNC in video-based music generation.
+These results highlight EMSYNC’s effectiveness in generating music that aligns with the input video both temporally and emotionally, across objective and subjective evaluations.
 By automatically generating soundtracks for user-provided videos, our work has the potential to streamline video production and enhance viewer engagement while offering a valuable framework for both machine learning researchers and multimedia content creators.
+Dataset EmoMV-C Ads
+Metric AVf AVs KL AVf AVs
+CMT 0.014 0.150 4.728 0.008 0.099
+Video2Music 0.010 0.245 9.789 0.002 0.107
+EMSYNC (Ours) 0.017 0.233 4.175 0.011 0.150
+Table 2: Objective results: AVf and AVs denote audio-video alignment in frame and second-levels respectively, where higher is better. KL denotes KL-divergence where lower is better.
+D M CMT Video2Music
+EMSYNC
+EmoMV-C
+MR 1.870 ± 0.102 2.317 ± 0.108 1.812 ± 0.112
+MQ 2.226 ± 0.099 1.952 ± 0.116 1.822 ± 0.113
+EM 2.111 ± 0.100 2.043 ± 0.115 1.846 ± 0.117
+TM 2.058 ± 0.098 2.067 ± 0.115 1.875 ± 0.120
+OM 2.106 ± 0.100 2.087 ± 0.114 1.808 ± 0.116
+Ads
+MR 1.949 ± 0.153 2.439 ± 0.155 1.612 ± 0.140
+MQ 2.469 ± 0.150 1.888 ± 0.144 1.643 ± 0.153
+EM 2.214 ± 0.153 2.051 ± 0.171 1.735 ± 0.154
+TM 2.245 ± 0.168 2.010 ± 0.159 1.745 ± 0.150
+OM 2.255 ± 0.153 2.122 ± 0.166 1.622 ± 0.145
+Table 3: Subjective results: Mean ranks ± 95% confidence margins for the subjective metrics, where 1 is the best possible and 3 is the worst possible ranking.
+The first two columns represent the dataset and metric, indicated by the headers D and M.
+The table is transposed for compactness.

### v3

@@
 Abstract
-We introduce EMSYNC, a video-based symbolic music generation model that aligns music with a video’s emotional content and temporal boundaries.
+Providing soundtracks for videos remains a costly and time-consuming challenge for multimedia content creators.
+We introduce EMSYNC, an automatic video-based symbolic music generator that creates music aligned with a video’s emotional content and temporal boundaries.
 It follows a two-stage framework, where a pretrained video emotion classifier extracts emotional features, and a conditional music generator produces MIDI sequences guided by both emotional and temporal cues.
 We introduce boundary offsets, a novel temporal conditioning mechanism that enables the model to anticipate upcoming video scene cuts and align generated musical chords with them.
-Unlike existing models, our approach retains event-based encoding, ensuring fine-grained timing control and expressive musical nuances.
-We also propose a mapping scheme to bridge the video emotion classifier, which produces discrete emotion categories, with the emotion-conditioned MIDI generator, which operates on continuous-valued valence-arousal inputs.
+We also propose a mapping scheme that bridges the discrete categorical outputs of the video emotion classifier with the continuous valence-arousal inputs required by the emotion-conditioned MIDI generator, enabling seamless integration of emotion information across different representations.
 Our method outperforms state-of-the-art models in objective and subjective evaluations across different video datasets, demonstrating its effectiveness in generating music aligned to video both emotionally and temporally.
-1 Introduction
-The online distribution of user-generated multimedia content is expanding at an exponential rate thanks to affordable, high-quality recording equipment and video editing software (Falkowski-Gilski and Uhl 2020).
-A key challenge in content creation is providing suitable soundtracks to enhance viewer engagement (Buhler 2018).
+Our demo and output samples are available at https://serkansulun.com/emsync.
+I Introduction
+The online distribution of user-generated multimedia content is expanding at an exponential rate thanks to affordable, high-quality recording equipment and video editing software [11].
+A key challenge in content creation is providing suitable soundtracks to enhance viewer engagement [4].
 However, unauthorized use of commercially published music infringes copyright, preventing monetization for creators on platforms like YouTube.
 Alternatives such as purchasing music, hiring composers, or searching for royalty-free tracks are often costly, time-consuming, or fail to ensure proper synchronization with video content.
-Automatic video-based music generation offers a promising solution to this problem (Di et al.
-2021; Zhuo et al.
-2023; Kang, Poria, and Herremans 2024).
-Existing approaches that generate music as audio waveforms lack editability (Lam et al.
-2024; Copet et al.
-2024).
-These methods compress all stages of music production, including composition, performance, editing, mixing, and mastering, into a single process, which limits creative control for musicians and producers.
-By contrast, generating music in a symbolic format like MIDI (Musical Instrument Digital Interface) offers greater flexibility.
-MIDI functions as a digital music score, encoding instrument names, note pitches, durations, and velocities.
-Musicians and producers can edit compositions, synthesize audio using digital audio workstations (DAWs), or record and refine performances.
+Automatic video-based music generation offers a promising solution to this problem [56, 8, 21, 46].
+Existing approaches generate music as audio waveforms, which lack editability [28, 6] because they collapse composition, performance, and mastering into a single, inseparable signal.
+By contrast, generating music in a symbolic MIDI format offers greater flexibility by maintaining these stages as discrete, modifiable events.
+MIDI functions as a digital score, encoding instrument names, note pitches, durations, and velocities.
+Because these elements are editable, musicians can modify compositions or refine performances in digital audio workstations (DAWs).
 This work focuses on generating music in MIDI format from arbitrary videos using deep neural networks (DNNs).
@@
 A major challenge in training video-to-MIDI DNNs is the absence of large-scale paired video-MIDI datasets.
-Existing datasets are either domain-specific, such as pianist hand videos (Koepke et al.
-2020), or contain limited samples, around 1k (Di et al.
-2021; Zhuo et al.
-2023).
-To address this, we develop a model that generates music for any video type by leveraging the Lakh MIDI Dataset (Raffel 2016), the largest available collection with 176,581 samples, to ensure diverse and high-quality outputs.
+Existing datasets are either domain-specific, such as pianist hand videos [26] or MIDI files with lyrics metadata [42], or contain limited samples, around 1k [8, 56].
+To address this, we develop a model that generates music for any video type by leveraging the Lakh MIDI Dataset [36], the largest available collection with 176,581 samples, to ensure diverse and high-quality outputs.
 Since the Lakh MIDI dataset lacks corresponding videos, we adopt a two-stage approach: extracting video features relevant to music generation and using them as conditioning inputs.
-Drawing from musicology literature, we identify two essential aspects for effective video-based music generation: emotional alignment (Deutsch 2007) and temporal synchronization (Prendergast 1992).
+Drawing from the musicology literature, we identify two essential aspects for effective video-based music generation: emotional alignment [7] and temporal synchronization [34].
 We name our model EMSYNC, as it aligns music with video by matching their emotions, and synchronizing their temporal boundaries.
-While our method is applicable to any selection of temporal boundary and emotion representation, we define musical boundaries using long-duration chords, video boundaries using scene cuts, and represent emotions with the valence-arousal model (Russell 1980).
-Specifically, we extract scene cut locations from the input video and guide the music generator to produce long-duration chords near those locations that are rhythmically and harmonically compatible with the rest of the generated music.
-We also use a pretrained video emotion classifier to estimate discrete emotion probabilities, map them to valence-arousal values, and condition our music generator accordingly.
+While our method is applicable to any selection of temporal boundary and emotion representation, we define musical boundaries using long-duration chords, video boundaries using scene cuts, and represent emotions with a valence-arousal model [38].
+Specifically, we extract scene cut locations from the input video and guide the music generator to produce long-duration chords near locations that are rhythmically and harmonically compatible with the rest of the generated music.
+We also use our pretrained video emotion classifier to estimate discrete emotion probabilities [44, 43], map them to valence-arousal values, and condition our music generator accordingly.
 Temporal conditioning in MIDI-generating models presents unique challenges.
-While deep transformer models can handle time-based data, such as videos, their sequence dimension correlates linearly with time due to a fixed frame rate (Arnab et al.
-2021; Liu et al.
-2022).
-In contrast, MIDI is typically processed using an event-based representation, where sequence and time dimensions are not linearly correlated (Oore et al.
-2020).
+While deep transformer models can handle time-based data, such as videos, their sequence dimension correlates linearly with time due to a fixed frame rate [1, 30].
+In contrast, MIDI is processed using an event-based representation in modern applications [8, 21, 46], where sequence and time dimensions are not linearly correlated [31].
 In event-based MIDI encoding, two primary token types exist: “note” and “time shift”.
@@
 Each time shift token specifies a time increment, forming a one-dimensional sequence where the position in the sequence does not directly correspond to the position in time.
-The key advantage of event-based encoding is the absence of a fixed time grid, allowing for subtle, expressive timing variations that reflect human musicianship.
-Unlike state-of-the-art video-based music generators that rely on fixed, coarse time grids (Di et al.
-2021; Kang, Poria, and Herremans 2024), we introduce a novel temporal conditioning method that preserves event-based encoding, enabling fine-grained temporal control.
+The key advantage of event-based encoding is the absence of a fixed time grid, allowing for expressive timing variations that reflect human musicianship.
+Unlike state-of-the-art video-based music generators that rely on fixed, coarse time grids [56, 8, 21, 46], we introduce a novel temporal conditioning method that preserves event-based encoding, enabling fine-grained temporal control.
 The selection of temporal features for video-music synchronization is also non-trivial.
-Existing work matches video motion speed and motion saliency with musical note density, producing a dense representation of temporal dynamics that leads to continuous conditioning and frequent changes in note density (Di et al.
-2021).
+Existing work matches video motion speed and motion saliency with musical note density, producing a dense representation of temporal dynamics that leads to continuous conditioning and frequent changes in note density [8, 46, 56, 21].
 Our inspection of the output samples from this approach identified two key limitations.
@@
 Second, because the mapping is dense, synchronization between music and video is difficult to perceive unless there are strong contrasts in motion speed and note density across consecutive sections.
-Motivated by works in musicology and soundtrack composition, we choose a temporal representation of video using sections, separated by sparse temporal boundaries of scene cuts (Cohen 1990; Prendergast 1992).
-To enable our music generator to anticipate future boundaries and align its output accordingly, we propose using boundary offsets.
+Motivated by works in musicology and soundtrack composition, we choose a temporal representation of video using sections, separated by the sparse temporal boundaries of scene cuts [5, 34].
+By focusing on these sparse boundaries rather than continuous video features, the model only needs to align with discrete temporal points instead of constantly modifying note density, therefore avoiding potential rhythmic instabilities.
+To achieve this, we introduce boundary offsets, which are auxiliary inputs for each token representing the time remaining until the next scene cut.
+This allows the autoregressive generator to anticipate upcoming boundaries at every step and align generated musical events with them, ensuring precise synchronization.
 We conduct objective and subjective evaluations to measure the temporal synchronization between input video and generated music, and demonstrate that our approach outperforms dense temporal conditioning.
-In addition, we tackle the challenge of using emotions as an intermediary to link music generation to video.
-Emotions are commonly represented with either categorical (Ekman 1971) or dimensional (continuous-valued) (Russell 1980) models.
-Based on prior findings from psychology and emotion research (Russell and Mehrabian 1977), we provide a mapping between categorical and dimensional emotion representations.
-This mapping enables the combined use of different emotion representations across data modalities, such as categorical emotions for video classification and dimensional emotions for MIDI generation, as in our case.
-This approach ensures flexibility and supports the use of training data labeled with either type of emotion representation.
-We compare our method to state-of-the-art video-based MIDI generators through objective and subjective evaluation on different datasets, where it outperforms the state-of-the-art methods.
-Our main contributions are as follows:
-  • We develop a video-based MIDI generator that surpasses state-of-the-art models across all subjective and a majority of the objective metrics for all datasets.
-  • We introduce boundary offsets for temporal conditioning in transformers, ensuring precise alignment between the input video and the output MIDI.
-  • Based on psychological research, we establish a mapping between discrete emotion categories and continuous valence-arousal values, an enabling emotional connection between the input video and the generated music.
-2 Related work
+In addition, we tackle the challenge of using emotions as an intermediary to link music generation with video.
+Our pretrained video emotion classifier outputs probabilities of categorical emotions [44], trained on Ekman‑6, one of the largest video-emotion datasets providing categorical labels [51].
+For music, we previously labeled the Lakh MIDI dataset [36] with valence-arousal values [41], producing a training set with a sample size 90 times that of the largest emotion-labeled MIDI dataset [32] and a total duration 187 times that of the longest [19].
+To bridge these different emotion representations, we introduce a mapping between categorical and dimensional emotions using results from prior user studies [37].
+This enables the use of larger datasets for improved performance, such as the Ekman-6 video dataset with categorical labels [51] and our previous MIDI dataset with dimensional labels [41], while supporting training on either type of emotion annotation and maintaining flexibility.
+We compare our method to state-of-the-art video-based MIDI generators through objective and subjective evaluation on different datasets. Our main contributions are as follows:
+  • We present a new framework to train a video-based MIDI generator that surpasses state-of-the-art models across all subjective and a majority of the objective metrics for all datasets.
+  • We introduce the concept of boundary offsets for temporal conditioning in transformers, ensuring precise alignment between the input video and the output MIDI.
+  • Based on psychological research [37], we propose a mapping between discrete emotion categories and continuous valence-arousal values, enabling the integration of multimodal data labeled with different emotion representations for video-to-music generation.
+II Related work
 While our method applies to arbitrary videos, several works focus on generating symbolic music for specific types of videos, such as those featuring human movements like dancing or instrumental performances.
-The Foley Music model (Gan et al.
-2020) generates MIDI from videos of musicians by processing body keypoint movements using a Graph Convolutional Network (Kipf and Welling 2017) and a Transformer (Vaswani et al.
-2017).
+The Foley Music model [12] generates MIDI from videos of musicians by processing body keypoint movements using a Graph Convolutional Network [25] and a Transformer [48].
 Similarly, Koepke et al.
-(2020) and Su et al.
-(2020) use deep neural networks with residual connections to generate symbolic music from videos of finger movements on piano keyboards.
-Due to their specialized nature, these approaches rely on datasets containing video-MIDI pairs, however, these datasets typically contain fewer than 1k samples (Zhao et al.
-2018; Li et al.
-2018; Koepke et al.
-2020; Su, Liu, and Shlizerman 2020).
-The RhythmicNet model employs a multi-stage process to generate music from dance videos by predicting beats and style, generating a drum track, and subsequently creating multitrack music (Su, Liu, and Shlizerman 2021).
+[26] and Su et al.
+[39] use deep neural networks with residual connections to generate symbolic music from videos of finger movements on piano keyboards.
+Due to their specialized nature, these approaches rely on datasets containing video-MIDI pairs, however, these datasets typically contain fewer than 1k samples [55, 29, 26, 39].
+The RhythmicNet model employs a multi-stage process to generate music from dance videos by predicting beats and style, generating a drum track, and subsequently creating multitrack music [40].
 Some studies explore the more general task of generating symbolic music for arbitrary videos.
-The most similar to our approach, the Controllable Music Transformer (CMT), generates music based on video features such as motion speed, motion saliency, and timing (Di et al.
-2021).
-CMT employs the Lakh Pianoroll Dataset (Dong et al.
-2018), the same dataset we use, and processes music using an extended Compound Word representation, where each token encodes type, beat/bar marking, note strength, note density, instrument, pitch, and duration (Hsiao et al.
-2021).
-While this representation reduces sequence length, it significantly increases input dimensionality.
-Consequently, CMT uses a low temporal resolution of four ticks per beat, translating to a time resolution of 125 milliseconds for a song at 120 beats per minute.
-Furthermore, note events align precisely with beat subdivisions, whereas human performance naturally deviates from beat subdivisions due to expressive timing (Fabian, Timmers, and Schubert 2014).
-In contrast, our model utilizes an event-based representation with independent time shift tokens at an 8 ms resolution.
-This resolution is sufficient to capture musical nuance while maintaining absolute time rather than rigid beat subdivisions (Oore et al.
-2020).
-Additionally, unlike CMT, our model integrates high-level emotional conditioning alongside low-level temporal features.
+The Controllable Music Transformer (CMT) generates music based on video features such as motion speed, motion saliency, and timing [8].
+It processes music using an extended Compound Word representation, where each token encodes type, beat/bar marking, note strength, note density, instrument, pitch, and duration [17].
 In a follow-up to CMT, Zhuo et al.
-introduced the Symbolic Music Videos dataset, containing video-MIDI pairs (2023).
-They sourced MIDI data from piano tutorial videos, automatically transcribing audio using the Onsets and Frames model, resulting in 1140 samples (Hawthorne et al.
-2018).
-Video features such as color histograms, RGB frame differences for motion, and high-level CLIP features (Radford et al.
-2021) were extracted to train three models sequentially: one generating chord sequences, another generating the melody, and a third generating the accompaniment.
-The Video2Music model shares similarities with our approach by utilizing both low-level video features and high-level emotional conditioning, but differs in application (Kang, Poria, and Herremans 2024).
+introduced the V-MusProd model alongside the Symbolic Music Videos dataset, containing video-MIDI pairs [56].
+They sourced MIDI data from piano tutorial videos, automatically transcribing audio using the Onsets and Frames model, resulting in 1140 samples [14].
+Video features such as color histograms, RGB frame differences for motion, and high-level CLIP features [35] were extracted to train three modules sequentially: one generating chord sequences, another generating the melody, and a third generating the accompaniment.
+The Video2Music model shares similarities with our approach by utilizing both low-level video features and high-level emotional conditioning, but differs in application [21].
 The authors compiled the MuVi-Sync dataset, consisting of 748 music videos labeled with musical attributes like note density, loudness, chords, and key.
-Their encoder-decoder transformer takes low- and high-level video features, along with user-provided primer chord and key, to generate chord sequences.
+Their encoder-decoder transformer takes low- and high-level video features to generate chord sequences.
 These sequences are then arpeggiated using fixed patterns to create the final MIDI output.
-However, this method’s reliance on a fixed time grid omits the subtle expressive timing found in human performances.
-Additionally, its use of fixed arpeggiation patterns limits musical diversity, and requiring user-defined chord and key inputs restricts accessibility for non-musicians.
-3 Methodology
+However, relying on fixed arpeggiation patterns limits the musical diversity of the generated pieces.
+The XMusic model generates MIDI for videos by analyzing the video in terms of emotional and temporal features [46].
+It uses Compound Word representation and performs dense, bar-by-bar conditioning.
+A “selector” module trained with human feedback chooses the best music from a batch of output samples in the end.
+The model is trained on a privately-collected closed-source dataset.
+CMT, V-MusProd, Video2Music, and XMusic all use a quantized temporal representation (32nd notes or coarser) and rely on dense temporal conditioning, often treating each musical bar independently.
+In contrast, our model adopts an event-based representation with independent time-shift tokens at an 8 ms resolution.
+This fine-grained resolution captures subtle timing variations, reflecting how humans naturally play music, rather than imposing rigid, quantized beat subdivisions [31].
+Furthermore, our use of sparse temporal conditioning in absolute time eliminates reliance on musical bars, enabling the model to handle any type of MIDI, including expressive performances that may not follow strict bar structures.
+III Methodology
 In this section, we present our video-based symbolic music generator.
-We match the output music to the input video based on emotions and temporal boundaries.
-Since there are no datasets where videos are paired with symbolic music, we employ a two-stage approach, i.e., use independent modules for video analysis and conditional music generation, and then combine them.
-Our pipeline is shown in Figure 1.
-We first highlight our music generator and how it is conditioned on emotions and temporal boundaries.
-We then describe how we incorporate temporal and emotional video features into our conditional music generator.
-3.1 Conditional music generator
-We extend an existing emotion-based MIDI generator trained with valence-arousal labels (Sulun, Davies, and Viana 2022) assigned to the Lakh MIDI dataset (Raffel 2016).
-By introducing temporal boundary conditioning and emotion mapping, we align music generation with input videos at inference time.
-Figure 2 illustrates the music generator, with the emotion-conditioning mechanism shown in the upper part and the boundary-conditioning mechanism in the lower part.
-Figure 1: Video-based music generation pipeline. The text and the image next to the arrows demonstrate sample values.
-Training dataset and preprocessing
-We train our music generator on the Lakh Pianoroll Dataset (LPD) (Dong et al.
-2018), which contains 174,154 pianorolls derived from the Lakh MIDI Dataset (Raffel 2016).
-We tokenize the pianorolls using an event-based symbolic music representation (Oore et al.
-2020).
+Since there are no datasets where videos are paired with symbolic music, we employ a two-stage approach, i.e., using independent modules for video analysis and conditional music generation, and then combine them.
+In order to make our contributions self-contained, we briefly summarize the existing components, namely, our video emotion classifier and emotion-based MIDI generator from prior work.
+Next, we describe how we integrate these components through a mapping scheme that bridges different emotion representation modalities, along with our novel temporal conditioning mechanism for generating MIDI that is temporally aligned with the input video.
+III-A Overview of prior components
+III-A1 Video emotion classification
+Our video emotion classifier exploits publicly available pretrained models for a multimodal video analysis and uses the resulting pretrained features to classify the emotions of arbitrary videos [44, 43].
+It employs pretrained models for automatic speech recognition (ASR), optical character recognition (OCR), facial expression classification, audio classification, and image understanding.
+Textual features obtained from ASR and OCR are further processed using a text sentiment classifier.
+Cross-attention layers [48] integrate and process these multimodal features, and a linear layer produces the final emotion probabilities.
+The model is trained on the Ekman-6 dataset [51], consisting of 1637 videos labeled with the six basic emotions derived from the original work of Ekman: anger, disgust, fear, joy, sadness, and surprise [10].
+III-A2 Emotion-based MIDI generation
+We have previously collected valence and arousal values for songs in the Lakh Pianoroll Dataset [9] and built a conditional transformer to use the resulting labeled dataset [41].
+To integrate continuous-valued valence and arousal features with discrete musical note tokens, we projected valence and arousal into a vector space using separate linear layers.
+Musical input tokens are projected into vectors of the same dimensionality using an embedding layer.
+These vectors are then concatenated with the projected valence and arousal vectors along the sequence dimension and fed into the transformer body with relative global attention [18].
+III-B Proposed method
+We extend our previous emotion-based MIDI generator [41] by introducing temporal boundary conditioning.
+Figure 1 illustrates the complete music generation architecture, where the upper section depicts the emotion-conditioning mechanism and the lower section shows the boundary-conditioning mechanism.
+Through the proposed emotion mapping, we integrate our video emotion classifier with the MIDI generator, forming a unified video-based MIDI generation framework, as shown in Figure 2.
+Figure 1: Our music generator.
+Numbers underneath valence and arousal are sample input values for illustration.
+Input tokens consist of discrete melodic events (e.g., ON, OFF, and TIMESHIFT) and special functional tokens (e.g., START, BAR, and CHORD).
+Boundary offsets are a sequence of scalar values, illustrated as the heights of the green bars.
+The plus sign indicates tensor addition.
+The dashed magenta boxes indicate concatenation.
+In sequence concatenation, the encoded valence and arousal vectors with shape (1, D) and the embedded input token sequence with shape (L, D) are concatenated along the first dimension, resulting in a tensor of shape (L+2, D).
+In feature concatenation, the encoded boundary offsets and positional embeddings, both with shape (L+2, D/2), are concatenated along the second dimension, producing a tensor of shape (L+2, D).
+Figure 2: Video-based music generation pipeline. The text and the image next to the arrows demonstrate sample values for illustration.
+III-B1 Training dataset and preprocessing
+We train our music generator on the Lakh Pianoroll Dataset (LPD) [9], which contains 174,154 pianorolls derived from the Lakh MIDI Dataset [36].
+We tokenize the pianorolls using an event-based symbolic music representation [31].
 Specifically, an ON (note on) token marks the start of a note, and an OFF (note off) token marks its end.
@@
 For example, a piano note with a MIDI pitch of 60 (C4) is denoted as PIANO_ON_60.
-Since a larger number of instruments increases vocabulary size, we use the Lakh Pianoroll Dataset-5 variant, where all instrument tracks are merged into five predefined categories: bass, drums, guitar, piano, and strings (Dong et al.
-2018).
+Since a larger number of instruments increases vocabulary size, we use the Lakh Pianoroll Dataset-5 variant, where all instrument tracks are merged into five predefined categories: bass, drums, guitar, piano, and strings [9].
 However, our method is adaptable to datasets with different instrument groupings.
@@
 We also use the START tokens to mark the beginning of the songs, the BAR tokens to indicate the musical bars, and the PAD tokens to standardize input sequence lengths in minibatches.
-We introduce the CHORD token to mark long-duration chords as temporal boundaries, as they often serve as anchor points in musical compositions (Bharucha and Krumhansl 1983).
-However, this is a design choice, and our method can accommodate any selection of temporal boundaries.
-To fully exploit available training data, we implement several strategies.
+We introduce the CHORD token to mark long-duration chords as temporal boundaries, as they often serve as anchor points in musical compositions [2].
+However, this is a design choice, and our method can accommodate any selection of temporal boundaries such as those separating melodic phrases or song sections.
+To fully exploit the available training data, we implement several strategies.
 Since we aim to generate multi-instrument compositions, we prioritize pieces with three or more instruments.
-However, one third of the Lakh MIDI dataset contains songs with only one or two instruments, with classical piano pieces dominating the single-instrument subset (Raffel 2016).
+However, one third of the Lakh MIDI dataset contains songs with only one or two instruments, with classical piano pieces dominating the single-instrument subset [36].
 As these samples still contain valuable information on musical structure, we avoid filtering them out.
@@
 If valence or arousal is unspecified, we assign it a NaN value and use a learned vector in place of its projection vector.
-Figure 2: Our music generator. Numbers underneath valence and arousal are sample values.
-Temporal boundary matching
+III-B2 Temporal boundary matching
 Our model generates music that is synchronized with the video’s temporal boundaries.
@@
 These chords are labeled by inserting a CHORD token before the first ON token of each chord.
-To allow the model to generate chords not only at video scene cut locations but also independently where musically appropriate, we randomly remove 20% of CHORD tokens during training.
 An example of a labeled chord is shown in the top row of Figure 3.
 In training, the temporal locations of chords from the ground-truth MIDI serve as input boundaries, whereas during video-based inference, we replace these boundaries with video scene cut locations.
-Chords are integral to a melody, providing harmonic and rhythmic support to surrounding notes, both preceding and following (Bharucha and Krumhansl 1983).
+Chords are integral to a melody, providing harmonic and rhythmic support to surrounding notes, both preceding and following [2].
 During inference, forcing a CHORD token into the sequence at a specific location may cause the chord to sound off-beat or overly abrupt.
@@
 These offsets are capped at a maximum value and, since our music generator is autoregressive, they are computed based on future chords rather than past ones.
-Furthermore, we amplify the loss of the CHORD token by a factor of 10, as it influences multiple preceding chord notes and plays a critical role in structuring the generated music.
 Sample offsets are illustrated in the bottom row of Figure 3.
-We process boundary offsets through a feed-forward network, generating boundary offset encodings, which are then concatenated with learned positional encodings (Wang et al.
-2020) along the feature dimension, as shown in the lower part of Figure 2.
+We process boundary offsets through a feed-forward network, generating boundary offset encodings, which are then concatenated with learned positional encodings [49] along the feature dimension, as shown in the lower part of Figure 1.
 This architectural choice is motivated by several key factors.
@@
 Second, we avoid adding boundary offset encodings directly to learned positional encodings to maintain a distinction between the two.
-Finally, recent studies suggest that decoder-only transformers can implicitly learn positional encodings through their internal weights, even without explicitly adding them (Haviv et al.
-2022; Kazemnejad et al.
-2024).
+Finally, recent studies suggest that decoder-only transformers can implicitly learn positional encodings through their internal weights, even without explicitly adding them [13, 22].
 Based on this insight, we halve the feature length of learned positional encodings and allocate the remaining feature space to boundary offset encodings.
 The resulting vector sequence consists of positional encodings augmented with boundary offset encodings.
-Following the standard transformer model, we add this sequence to the token embeddings (Vaswani et al.
-2017) before passing it to the transformer body with relative global attention (Huang et al.
-2018).
+Following the standard transformer model, we add this sequence to the token embeddings [48] before passing it to the transformer body with relative global attention [18].
 The transformer’s input is represented as:
-$\begin{split}\mathbf{X}_{\text{trans}}&=\text{Concat}_{s}\left(\text{FFN}_{v}(x_{v}),\text{FFN}_{a}(x_{a}),\text{Embedding}(\mathbf{x}_{t})\right)\\ &+\text{Concat}_{f}(\text{FFN}_{b}(\mathbf{b})+\mathbf{W_{pe}})\end{split}$
-where $x_{v}$ and $x_{a}$ are the valence and arousal inputs, $\mathbf{x}_{t}$ is the input token sequence, $\mathbf{b}$ represents the input boundary offsets, and $\mathbf{W_{pe}}$ is the learned positional encoding.
+$\begin{split}\mathbf{X}_{\text{trans}}&=[\text{FFN}_{v}(x_{v}),\,\text{FFN}_{a}(x_{a}),\,\text{Embedding}(\mathbf{x}_{t})]_{1}\\ &\quad+[\text{FFN}_{b}(\mathbf{b}),\,\mathbf{W}_{\text{pe}}]_{2}\end{split}$
+$\text{FFN}_{v}(x_{v}),\text{FFN}_{a}(x_{a})\in\mathbb{R}^{1\times D},\quad\text{Embedding}(\mathbf{x}_{t})\in\mathbb{R}^{L\times D},$
+$\text{FFN}_{b}(\mathbf{b}),\mathbf{W}_{\text{pe}}\in\mathbb{R}^{(L+2)\times\frac{D}{2}},\quad\;\;\,\mathbf{X}_{\text{trans}}\in\mathbb{R}^{(L+2)\times D}$
+where $x_{v}$ and $x_{a}$ are the valence and arousal inputs, $\mathbf{x}_{t}$ is the input token sequence, $\mathbf{b}$ represents the boundary offsets, and $\mathbf{W}_{\text{pe}}$ is the learned positional encoding.
 The feed-forward networks $\text{FFN}_{v}$, $\text{FFN}_{a}$, and $\text{FFN}_{b}$ process the valence, arousal, and boundary offset inputs, respectively.
-The operations $\text{Concat}_{s}$ and $\text{Concat}_{f}$ denote concatenation along the sequence and feature dimensions, respectively.
-Figure 3: Graphical illustration of boundaries.
+The plus sign ++ denotes tensor addition.
+The first concatenation $\cdot\,]_{1}$ is along the sequence (first) dimension producing a sequence of length $L+2$, and the second concatenation $\cdot\,]_{2}$ is along the feature (second) dimension producing a feature of size $D$.
+Figure 3: Illustration of boundaries.
 Top: In symbolic music, a chord with three or more simultaneous notes and a duration exceeding a set threshold defines a musical boundary and is used during training.
+As CHORD is a special token without duration or pitch, it is omitted from the pianoroll representation.
+Its temporal location, however, aligns with the red line indicating the boundary.
 Middle: In video, scene cuts serve as video boundaries and are used during inference.
 Bottom: Boundary offset represents the temporal distance to the next boundary.
+Each offset is a scalar value, visualized as the height of the corresponding green bar.
 These figures are illustrative; offsets do not perfectly align with the music except at the boundary.
@@
 If the model generates a CHORD token, we calculate the absolute difference between the time cursor and each input boundary.
-The boundary with a difference smaller than the predefined sensitivity threshold of 1 s is considered successfully generated.
+The boundary with a difference smaller than the predefined sensitivity threshold $\xi$ is considered successfully generated.
 We then remove these boundaries from future offset calculations by replacing them with infinity.
 Next, we compute the boundary offset for the generated token, regardless of its type.
-This offset is determined as the distance to the next closest boundary, capped by a maximum offset value.
+This offset represents the distance to the next closest boundary.
+We compute it by subtracting the time cursor from each input boundary and replacing any negative values (corresponding to past boundaries) with infinity to ignore them.
+We then take the minimum of the resulting values as the offset to the nearest boundary and clamp it to a maximum value if necessary.
+The resulting boundary offset and generated token are appended to their respective lists to be used in the next timestep.
 For simplicity, Algorithm 1 is presented for a single sample, but in practice, this operation is performed in minibatches.
@@
 Music generation continues until the time cursor reaches the duration of the video.
-After synthesizing the generated MIDI into an audio waveform, we apply a 3-second fade-out effect to prevent an abrupt ending and ensure the music matches the duration of the input video.
 Algorithm 1 Creating boundary offsets during music generation in inference.
 Input: List of input boundaries (in seconds), $\mathbf{b}$; video duration, $d$; valence, $x_{v}$; arousal $x_{a}$
-Parameter: Sensitivity, $\xi$; maximum offset, $\delta_{max}$; generated boundary mask, $\mathbf{m_{b}}$; token type $t_{t}$; token value, $t_{v}$; time cursor, $c$; generation function including forward-pass and sampling, g
+Parameter: Sensitivity, $\xi$; distances to all boundaries, $\boldsymbol{\delta}_{b}$; maximum offset, $\delta_{\max}$; generated boundary mask, $\mathbf{m_{b}}$; token type $t_{t}$; token value, $t_{v}$; time cursor, $c$; generation function including forward-pass and sampling, g
 Output: List of generated tokens $\mathbf{t}$; list of boundary offsets, $\boldsymbol{\delta}$
@@
   else if $t_{t}$ == CHORD then
-   $\mathbf{m_{b}}=\left|t-\mathbf{b}\right|<\xi$.
+   $\mathbf{m_{b}}=\left|c-\mathbf{b}\right|<\xi$.
    $\mathbf{b}\left[\mathbf{m_{b}}\right]=+\infty$.
   end if
+  $\boldsymbol{\delta}_{b}=\mathbf{b}-c$
+  $\boldsymbol{\delta}_{b}[\boldsymbol{\delta}_{b}<0]=+\infty$
 […]
@@
 end while
-3.2 Scene detection
+III-B3 Scene detection
 During video-based inference, we first extract the video’s temporal boundaries, specifically its scene cut locations.
-As scene detection is a solved problem, we employ the off-the-shelf FFmpeg software [^1] https://www.ffmpeg.org/.
+As scene detection is a well-studied problem with established solutions, we employ the off-the-shelf FFmpeg software [^1] https://www.ffmpeg.org/.
 As we generate musical chords near scene boundaries, we aim to avoid an excessive number of densely spaced chords.
-We therefore apply a difference filter to the extracted scene cuts and remove those occurring less than 4 s apart.
+We therefore apply a difference filter to the extracted scene cuts and remove those occurring less than $t_{\text{diff}}$ seconds apart.
 The remaining timestamps serve as the input boundaries for the music generator.
-3.3 Emotion mapping
-We use a pretrained emotion classifier to extract emotions from input videos as probability distributions over discrete emotion categories (Sulun, Viana, and Davies 2024).
+III-B4 Emotion mapping
+We use our pretrained emotion classifier [44] to extract emotions from input videos as probability distributions over Ekman’s basic emotion categories [10].
 Our music generator is conditioned on emotions represented as valence and arousal values ranging from -1 to 1, rather than discrete categories.
-Since the emotion classifier outputs discrete emotion categories, we first map them to the valence-arousal plane (Russell 1980) before feeding them into the music generator.
-Russell and Mehrabian conducted user studies to find the corresponding valence and arousal values of discrete emotion categories (1977).
+Since the emotion classifier outputs discrete emotion categories, we first map them to the valence-arousal plane [38] before feeding them into the music generator.
+Russell and Mehrabian conducted user studies to find the corresponding valence and arousal values of discrete emotion categories [37].
 They presented their findings in a table that contains the mean and standard deviation values of valence and arousal for each categorical emotion.
-We extracted and used the values corresponding to Ekman’s six basic emotions (1971), which are presented in Table 1.
+We extracted and used the values corresponding to Ekman’s six basic emotions [10], which are presented in Table I.
+While Russell and Mehrabian presented one of the oldest user studies that examine the relationship between dimensional and categorical emotion representations, their study remains the most comprehensive in terms of number of users ($N=300$) and number of categorical emotions ($K=151$), and is still employed in recent machine learning applications [27].
+Similar and newer user studies from Hoffmann et al.
+($N=70$, $K=22$) [15] and Trnka et al.
+($N=187$, $K=16$) [47] are more limited, and do not include all of Ekman’s basic emotions.
+Using the output probabilities of the video classifier along with the means and standard deviations of each emotion, we construct a mixture of Gaussian distributions.
+Valence-arousal values can be obtained either by sampling an emotion category based on the output probabilities and then drawing from its corresponding Gaussian, or by using the mixture mean computed as the weighted average of the category means.
+Additionally, we introduce a parameter $\mu_{\max}$ representing the maximum absolute value of means across all emotion categories.
+Using $\mu_{\max}$, we compute a single scaling coefficient applied uniformly to all valence-arousal distributions.
+This approach allows users to adjust the conditioning of the music generator, choosing between a wider or narrower range of emotions.
+Although not used in our model, we also formalize a method to invert this mapping, that is, to map valence-arousal values to emotion categories by comparing the valence-arousal coordinates to each category’s distribution defined by its mean and standard deviation.
+This involves calculating a distance metric, such as Euclidean or Mahalanobis distance, or computing the likelihood assuming Gaussian distributions, and then assigning the category with the smallest distance or highest likelihood.
+This provides a generalizable framework for integrating dimensional and categorical emotion representations in multimodal systems.
+TABLE I: Valence and arousal means and standard deviations for Ekman’s six basic emotions [10], derived from Russell and Mehrabian’s user study [37].
 Valence Arousal
@@
 surprise -0.40 0.30 -0.67 0.27
-Table 1: Means and standard deviations of valence and arousal values corresponding to Ekman’s categorical emotions (1971), obtained from the user studies of Russell and Mehrabian (1977).
-Using the output probabilities of the video classifier along with the means and standard deviations of each emotion, we construct a mixture of Gaussian distributions.
-Valence-arousal values can be obtained either by sampling an emotion category based on the output probabilities and then drawing from its corresponding Gaussian, or by using the mixture mean computed as the weighted average of the category means.
-Additionally, we introduce a parameter that represents the maximum absolute value of means across all emotion categories.
-Using this parameter, we compute a single scaling coefficient applied to all valence-arousal distributions.
-This approach allows users to adjust the conditioning of the music generator, choosing between a wider or narrower range of emotions.
-Although not used in our work, it is possible to invert this mapping, that is, to map valence-arousal values to emotion categories by comparing the valence-arousal coordinates to each category’s distribution defined by its mean and standard deviation.
-This involves calculating a distance metric, such as Euclidean or Mahalanobis distance, or computing the likelihood assuming Gaussian distributions, and then assigning the category with the smallest distance or highest likelihood.
-4 Experimental setup
-We perform objective and subjective evaluations of our overall model using two datasets, against the open-source state-of-the-art models that generate MIDI from arbitrary videos: Video2Music (Kang, Poria, and Herremans 2024) and CMT (Di et al.
-2021).
-4.1 Implementation details
+IV Experimental setup
+We perform objective and subjective evaluations of our overall model using two datasets, against open-source state-of-the-art models that generate MIDI from arbitrary videos.
+Comparison with V-MusProd [56] and XMusic [46] was not possible, as their code is not publicly available.
+We therefore compare our method against Video2Music [21] and CMT [8].
+IV-A Implementation details
 Our music generator consists of 11 layers, 8 attention heads, and a model dimensionality of 512.
-To ensure a fair comparison, we align our model size with the compared models: our model has 37M parameters, compared to 39M in CMT (Di et al.
-2021) and 33M in Video2Music (Kang, Poria, and Herremans 2024).
+To ensure a fair comparison, we align our model size with the compared models: our model has 37M parameters, compared to 39M in CMT [8] and 33M in Video2Music [21].
 Training is done with a context length of 1216 and a batch size of 64.
-The model is implemented in PyTorch (Paszke et al.
-2019) and trained on a single NVIDIA A100 80GB GPU.
-For video-based inference in evaluation, we set the maximum absolute value of the emotion category means to 0.8.
+The model is implemented in PyTorch [33] and trained on a single NVIDIA A100 80GB GPU.
+We train our model using cross-entropy loss with a learning rate of 2e-4 for the first 300k steps, followed by 5e-5 for the next 300k steps.
+We use the Adam optimizer with gradient clipping to a norm of 1 [24].
+For data augmentation, we transpose the pitches of all instruments, except drums, by a randomly chosen number of semitones between -3 and 3, inclusive.
+To prevent the model from becoming overly reliant on the CHORD token for chord generation, we randomly remove 20% of these tokens during training.
+Furthermore, we amplify the loss of the CHORD token by a factor of 10, as it influences multiple preceding chord notes and plays a critical role in structuring the generated music.
+We set the maximum boundary offset $\delta_{\max}$ to 8 seconds.
+We reserve 5% of the training data for validation.
+These training hyperparameters are selected via grid search to minimize validation loss.
+For video-based inference in evaluation, we set the maximum absolute value of the emotion category means $\mu_{\max}$ to 0.8.
+We obtain valence–arousal values as the mean of the weighted mixture of Gaussians derived from the predicted emotion categories.
+Although this conditioning is deterministic, the music generator remains stochastic due to probabilistic token sampling during generation.
+The difference filter discards scene timestamps that are less than $t_{\text{diff}}=4$ seconds apart.
+We set the boundary sensitivity threshold $\xi$ to 1 second.
 To ensure consistency, all methods generate MIDI outputs that are synthesized into audio waveforms using Fluidsynth [^2] https://www.fluidsynth.com, followed by peak normalization to -3 dB.
+We apply a 3-second fade-out effect to prevent an abrupt ending and ensure the music matches the duration of the input video.
 The audio is then merged with the input video by overwriting its audio channel.
 For all methods, outputs are generated in a single run across all datasets using a fixed random seed.
-We train our model using cross-entropy loss with a learning rate of 2e-4 for the first 300k steps, followed by 5e-5 for the next 300k steps.
-We use the Adam optimizer with gradient clipping to a norm of 1 (Kingma and Ba 2015).
-For data augmentation, we transpose the pitches of all instruments, except drums, by a randomly chosen number of semitones between -3 and 3, inclusive.
-We reserve 5% of the training data for validation.
-Hyperparameters are selected via grid search to minimize validation loss.
-4.2 Evaluation datasets
-For objective and subjective evaluation, we first use the EmoMV-C dataset, which consists of music videos containing music audio tracks to be used as ground truth for objective evaluation (Thao, Roig, and Herremans 2023).
+The inference hyperparameters are chosen by manually inspecting outputs on videos not included in the evaluation.
+IV-B Evaluation datasets
+For objective and subjective evaluation, we first use the EmoMV-C dataset, which consists of music videos containing music audio tracks to be used as ground truth for objective evaluation [45].
 We exclude auditory features during the emotion classification to prevent data leakage.
-To assess the model’s generalization across different video types, we also include the Pittsburgh Advertisements (Ads) dataset, which contains advertisement videos that may or may not feature music (Hussain et al.
-2017).
-As advertisements often rely on music to maximize viewer engagement, the Ads dataset provides a real-world scenario for evaluation.
-(Zander 2006).
+To assess the model’s generalization across different video types, we also include the Pittsburgh Advertisements (Ads) dataset, which contains advertisement videos that may or may not feature music [20].
+As advertisements often rely on music to maximize viewer engagement, the Ads dataset provides a real-world scenario for evaluation [54].
 We use the full EmoMV-C validation split, which contains 48 thirty-second videos.
 We filter the Ads dataset to match the total duration of EmoMV-C.
-We select videos associated with the four basic emotions commonly used in music emotion classification: cheerful, calm, angry, and sad (Yang et al.
-2008).
-We then remove videos shorter than 1 minute.
+We select videos associated with the four basic emotions commonly used in music emotion classification [52], covering the quadrants of Russell’s valence-arousal model [38]: cheerful, calm, angry, and sad, providing a broad coverage of the emotional space for evaluation.
+While some of the chosen emotions do overlap with Ekman’s categories [10], we avoid explicitly using Ekman’s emotions to select the evaluation videos.
+This prevents biasing the evaluation toward our method, since our video emotion classifier is trained to output Ekman’s categories.
+We exclude videos shorter than 1 minute.
 To ensure an unbiased selection, we use YouTube IDs, which are generated randomly by YouTube.
@@
 The resulting evaluation videos are then fed into the compared models.
-4.3 Objective evaluation
+Figure 4: Sample view of the survey used in the subjective evaluation. Participants were asked to drag the method boxes from the left onto the ranking boxes on the right.
+IV-C Objective evaluation
 We first evaluate music similarity using the ground-truth audio from the EmoMV-C dataset.
 Although the ground truth is not in MIDI, all methods generate MIDI outputs that are synthesized to audio, enabling a fair comparison in the audio domain.
-We evaluate the Kullback-Leibler Divergence (KL) between the labels of generated and ground-truth music, calculated using a pretrained music tagging model (Won et al.
-2020).
+We evaluate the Kullback-Leibler Divergence (KL) between the labels of generated and ground-truth music, calculated using a pretrained music tagging model [50].
 This measures how closely the distribution of semantic tags in the generated music matches that of the ground-truth.
-We avoid using Fréchet Audio Distance, as it compares overall audio distributions suited for unconditional generation (Kilgour et al.
-2019), whereas video-based music generation requires per-sample evaluation to assess alignment with each video’s unique content.
-We additionally evaluate temporal audio-video alignment (AV-Align) score, which is the Intersection-over-Union between video and audio peaks (Yariv et al.
-2024), using both EmoMV-C and Ads datasets.
-Video peaks are detected from the mean optical flow magnitude (Horn and Schunck 1981), while audio peaks correspond to audio onsets (Böck and Widmer 2013).
-We report AV-Align scores using windows of one video frame (33.4 ms at 30 FPS) and one second.
-Each window is applied to detect local peaks in both video and audio and then used to match peaks within the same temporal range for alignment.
-4.4 Subjective evaluation
+We avoid the commonly used Fréchet Audio Distance, as it compares overall audio distributions suited for unconditional generation [23], whereas video-based music generation requires per-sample evaluation to assess alignment with each video’s unique content.
+We additionally evaluate temporal audio-video alignment (AV-Align) score, which is the Intersection-over-Union between video and audio peaks [53], using both EmoMV-C and Ads datasets.
+Video peaks are detected from the mean optical flow magnitude [16], while audio peaks correspond to audio onsets [3].
+We calculate AV-Align scores at two temporal resolutions (windows): one video frame (33.3 ms at 30 FPS) and one second.
+For each window, local peaks in both video and audio are detected and matched within the same temporal range to assess alignment.
+As an internal evaluation, we additionally report the alignment between the generated MIDI chords (specifically the CHORD token) and video scene cuts using the same 33.3 ms and 1 s windows.
+As baseline methods lack an explicit boundary-matching architecture, these alignment scores are reported exclusively for EMSYNC to validate the effectiveness of its temporal conditioning mechanism.
+Finally, we report the average runtime for each model when generating music for a one-minute video, including model initialization, video analysis, and music generation.
+IV-D Subjective evaluation
 Using the evaluation videos from the EmoMV-C and the Ads datasets, we generate accompanying music with the three models and create survey pages, each containing two videos.
-For each video, the three music versions generated by the compared models are presented side by side with anonymized model names, with each model’s output appearing equally in the left, center, and right positions.
+For each video, the three music versions generated by the compared models are presented side by side with anonymized model names, with each model’s output appearing an equal number of times in the left, center, and right positions.
 We also provide output samples from eight videos online, with model names deanonymized and each model’s output appearing in the same position for clarity.
-[^3] https://em-sync.github.io
-We enrolled 153 participants and asked them to rank the three models based on the standard criteria used in previous works (Di et al.
-2021; Kang, Poria, and Herremans 2024): Music Richness (MR) — the richness and diversity of the music, independent of the video content; Music Quality (MQ) — the overall quality of the music, independent of the video content; Emotion Match (EM) — how well the music matches the video in terms of emotion; Timing Match (TM) — how well the music synchronizes with the video in terms of rhythm and timing; and Overall Match (OM) — how well the music matches the video as a whole.
-Each model is assigned a unique ranking of 1 (best), 2, or 3 (worst).
-5 Results and conclusion
-Table 2 presents objective evaluation results for EMSYNC, Video2Music (Kang, Poria, and Herremans 2024), and CMT (Di et al.
-2021), on the EmoMV-C and Ads datasets.
+[^3] https://serkansulun.com/emsync
+We enrolled 153 remote participants who provided informed consent before taking part in the survey.
+They were asked to rank the three models based on the standard criteria used in previous works [8, 21]:
+  • Music Richness (MR) — the richness and diversity of the music, independent of the video content.
+  • Music Quality (MQ) — the overall quality of the music, independent of the video content.
+  • Emotion Match (EM) — how well the music matches the video in terms of emotion.
+  • Timing Match (TM) — how well the music synchronizes with the video in terms of rhythm and timing.
+  • Overall Match (OM) — how well the music matches the video as a whole.
+Each model is assigned a unique ranking of 1 (best), 2, or 3 (worst). Figure 4 shows a sample view of the user survey.
+V Results and conclusion
+TABLE II: Objective comparison with the state-of-the-art
+Dataset EmoMV-C Ads
+Metric AVf AVs KL AVf AVs RT
+CMT 0.014 0.150 4.728 0.008 0.099 3.55
+Video2Music 0.010 0.245 9.789 0.002 0.107 1.61
+EMSYNC (Ours) 0.017 0.233 4.175 0.011 0.150 1.42
+Table II presents objective evaluation results for EMSYNC, Video2Music [21], and CMT [8], on the EmoMV-C and Ads datasets.
+AVf and AVs denote audio-video alignment in frame and second-levels, respectively, where higher is better.
+KL denotes the KL-divergence, where lower is better.
+RT denotes average runtime in minutes to generate music for a 1-minute video, including model initialization, video analysis, and music generation, where lower is better.
 EMSYNC outperforms both baselines across a majority of the metrics.
@@
 We don’t calculate KL divergence because the Ads dataset doesn’t include ground truth music.
-Table 3 demonstrates the subjective results in terms of average user rankings with 95% confidence margins.
-A lower average score represents a better ranking, with 1 being the best and 3 the worst possible.
-EMSYNC achieves the best rankings across all metrics and datasets.
-The difference is more pronounced in video-music correspondence metrics such as emotion match and timing match, highlighting our model’s stronger alignment to video.
-For overall match, our method outperforms the others with non-overlapping 95% confidence intervals, reflecting a decisive user preference for EMSYNC in video-based music generation.
-These results highlight EMSYNC’s effectiveness in generating music that aligns with the input video both temporally and emotionally, across objective and subjective evaluations.
-By automatically generating soundtracks for user-provided videos, our work has the potential to streamline video production and enhance viewer engagement while offering a valuable framework for both machine learning researchers and multimedia content creators.
-Dataset EmoMV-C Ads
-Metric AVf AVs KL AVf AVs
-CMT 0.014 0.150 4.728 0.008 0.099
-Video2Music 0.010 0.245 9.789 0.002 0.107
-EMSYNC (Ours) 0.017 0.233 4.175 0.011 0.150
-Table 2: Objective results: AVf and AVs denote audio-video alignment in frame and second-levels respectively, where higher is better. KL denotes KL-divergence where lower is better.
-D M CMT Video2Music
-EMSYNC
-EmoMV-C
-MR 1.870 ± 0.102 2.317 ± 0.108 1.812 ± 0.112
+EMSYNC also achieves the fastest runtime among the evaluated methods.
+For the EmoMV-C and Ads datasets, EMSYNC achieves chord-scene cut alignment scores of 0.297 and 0.317 at the 33.3 ms resolution, and 0.899 and 0.909 within a one-second window, respectively, demonstrating its capabilities in aligning temporal boundaries through its temporal conditioning mechanism.
+TABLE III: Subjective comparison with the state-of-the-art:
+Mean ranks ± 95% confidence margins
+Dataset Metric CMT Video2Music EMSYNC (Ours)
+EmoMV-C MR 1.870 ± 0.102 2.317 ± 0.108 1.812 ± 0.112
 MQ 2.226 ± 0.099 1.952 ± 0.116 1.822 ± 0.113
@@
 OM 2.106 ± 0.100 2.087 ± 0.114 1.808 ± 0.116
-Ads
-MR 1.949 ± 0.153 2.439 ± 0.155 1.612 ± 0.140
+Ads MR 1.949 ± 0.153 2.439 ± 0.155 1.612 ± 0.140
 MQ 2.469 ± 0.150 1.888 ± 0.144 1.643 ± 0.153
@@
 OM 2.255 ± 0.153 2.122 ± 0.166 1.622 ± 0.145
-Table 3: Subjective results: Mean ranks ± 95% confidence margins for the subjective metrics, where 1 is the best possible and 3 is the worst possible ranking.
-The first two columns represent the dataset and metric, indicated by the headers D and M.
-The table is transposed for compactness.
+Table III demonstrates the subjective results in terms of average user rankings with 95% confidence margins.
+A lower average score represents a better ranking, with 1 being the best and 3 the worst possible.
+EMSYNC achieves the best average ranking across all metrics and datasets.
+The difference is more pronounced in video-music correspondence metrics such as emotion match and timing match, highlighting our model’s stronger alignment to video.
+For the overall match, our method outperforms the others with non-overlapping 95% confidence intervals, reflecting a decisive user preference for EMSYNC in video-based music generation.
+Figure 5 further shows the distribution of rankings assigned by participants in the subjective evaluation.
+EMSYNC achieves the highest frequency of the top rank (Rank 1) across all metrics and datasets.
+Figure 5: Distributions of rankings assigned by the participants in the subjective evaluation using EmoMV-C (top) and Ads (bottom) datasets.
+These results highlight EMSYNC’s effectiveness in generating music that aligns with the input video both temporally and emotionally, across objective and subjective evaluations.
+By automatically generating soundtracks for user-provided videos, our work has the potential to streamline video production and enhance viewer engagement while offering a valuable framework for both machine learning researchers and multimedia content creators.
+In future work, we aim to incorporate more complex musical boundaries, such as those that separate melodic segments or motifs.
+We will also enable dynamic emotional conditioning to adapt the music continuously to evolving complex emotional content in videos.
+VI Acknowledgments
+This work is co-financed by Component 5 - Capitalization and Business Innovation, integrated in the Resilience Dimension of the Recovery and Resilience Plan within the scope of the Recovery and Resilience Mechanism (MRR) of the European Union (EU), framed in the Next Generation EU, for the period 2021 - 2026, within project NEXUS, with reference 53, and the fellowship from FCT - Fundação para a Ciência e a Tecnologia (2022.09594.BD).